### v1

# HydroGEM: A Self Supervised Zero Shot Hybrid TCN Transformer Foundation Model for Continental Scale Streamflow Quality Control

Abstract
Real-time streamflow monitoring networks generate millions of observations annually, yet maintaining data quality across thousands of remote sensors remains labor-intensive and difficult to scale. We introduce HydroGEM (Hydrological Generalizable Encoder for Monitoring), a foundation model for continental-scale streamflow quality control. HydroGEM uses a two-stage training approach: self-supervised pretraining on 6.03 million clean sequences from 3,724 USGS stations learns general hydrological representations, followed by fine-tuning with synthetic anomalies injected for detection and reconstruction. A hybrid Temporal Convolutional Network(TCN)-Transformer architecture (containing 14.2M parameters) captures both local temporal patterns and long-range dependencies, while hierarchical normalization handles learning across six orders of magnitude in discharge. On a held-out synthetic test set comprising 799 stations where 18 expert-validated anomaly types are observed, HydroGEM achieves F1 = 0.792 for detection accuracy and 68.7% reconstruction-error reduction, outperforming the strongest baseline by 36.3%. Zero-shot transfer to 100 Environment and Climate Change Canada (ECCC) stations—unseen during training—yields F1 = 0.586, exceeding all baselines and demonstrating cross-national generalization. The model maintains consistent detection across correction magnitudes (1–100%) and aligns with operational seasonal patterns, with peak flagging-rates during winter ice-affected periods matching hydrologists’ correction behavior. Architectural separation between simplified training anomalies and complex physical-space test anomalies confirms that performance reflects learned hydrometric principles rather than pattern memorization. HydroGEM is designed for human-in-the-loop workflows—outputs are quality control suggestions requiring expert review, not autonomous corrections. By automating routine screening while preserving human oversight, HydroGEM offers a scalable approach to expanding hydrologists’ effective reach across growing monitoring networks.
keywords:
Streamflow Quality Control , Foundation Model , Anomaly Detection , Self-Supervised Learning , Hydrological Monitoring , Deep Learning , Zero-Shot Transfer
journal: Environmental Modelling and Software
[…]
{highlights}
Continental-scale foundation model trained on 3724 USGS sites with 6.03 million sequences
Two-stage self-supervised pretraining with synthetic anomaly injection for detection
Achieves F1 = 0.792 detection accuracy and 68.7% reconstruction-error reduction
Zero-shot transfer to Canadian ECCC stations shows cross-national generalization
Human-in-the-loop design for operational quality control workflows
1 Introduction
Real-time hydrological monitoring networks provide essential observations for water resources management, flood forecasting, ecosystem protection, and climate change adaptation [kratzert2019ungauged, kratzert2022caravan, sterle2024camelschem]. The United States Geological Survey operates more than 10,000 stream gauging stations that collectively produce millions of paired discharge and stage measurements each month, forming the backbone of decision-making across federal, state, and local agencies [kratzert2019towards, sterle2024camelschem]. The value of these observations depends directly on data quality [nearing2022lstm].
However, maintaining quality assurance across thousands of remote sensors in harsh environments is difficult. Sensor drift, equipment faults, ice effects, rating curve changes, clock errors, and transmission problems frequently degrade data [sun2019review, dogo2019survey]. Current operational quality control relies primarily on expert hydrologists who visually inspect time series, reason about discharge-stage relationships, and apply site-specific judgment [nearing2022lstm, goldstein2015peeking]. While this expert-centered workflow delivers high-quality data, it does not scale with expanding networks and rising temporal resolution. The bottleneck is not lack of expertise—it is the mismatch between growing data volumes and limited expert capacity.
Machine learning approaches offer potential solutions but face two fundamental obstacles at continental scale. First, hydrological systems are extremely heterogeneous: discharge spans six orders of magnitude across sites, and flow generation mechanisms vary with climate, physiography, and human modification [kratzert2019towards, kratzert2022caravan, lees2021hydrological]. Models trained on individual sites or small regional datasets fail to generalize to unseen locations. Second, labeled anomaly datasets are scarce. Agency quality flags are not designed as ground truth labels that cleanly separate sensor faults from physical phenomena, and comprehensive labeled datasets across thousands of sites do not exist [dogo2019survey, belay2023unsupervised].
Foundation models trained with self-supervised objectives on massive unlabeled datasets provide a path forward [devlin2019bert, dooley2023forecastpfn]. In weather and climate science, transformers pretrained on reanalysis data now match or exceed numerical models for medium-range forecasting [nguyen2023climax, pathak2022fourcastnet, lam2023graphcast]. In hydrology, multi-basin training has enabled transfer learning to ungauged catchments [kratzert2019towards, kratzert2022caravan]. Yet foundation model principles have not been systematically applied to quality control for in situ sensor networks, where challenges include extreme scale heterogeneity, irregular multi-site time series, real-time processing requirements, and mandatory human oversight.
HydroGEM.
We introduce HydroGEM (Hydrological Generalizable Encoder for Monitoring), a self-supervised foundation model for continental-scale streamflow quality control. HydroGEM qualifies as a foundation model because it: (1) trains on massive diverse data (3724 sites spanning six orders of magnitude in discharge), (2) uses self-supervised pretraining to learn general hydrological representations, (3) shows zero-shot transfer to unseen sites and countries, and (4) uses a modular architecture that allows task-specific adaptation.
HydroGEM uses a two-stage training approach. Stage 1 pretrains a hybrid TCN-Transformer backbone on 6.03 million clean sequences from 3,724 USGS sites using masked reconstruction. A hierarchical normalization scheme—combining log transforms, site-specific standardization, and explicit scale embeddings—allows learning across extreme magnitude ranges while preserving physically meaningful scale structure [goodfellow2016deep]. Stage 2 fine-tunes the pretrained backbone with a detection head using on-the-fly synthetic anomaly injection rather than curated labels [dooley2023forecastpfn, kim2024enhancing]. The training injector applies simplified corruptions (drift, spikes, flatlines, dropouts, clock shifts) in normalized space to encourage learning of fundamental hydrometric consistency principles rather than memorization of specific anomaly signatures. Clean data preservation exceeds 97%, ensuring the model does not corrupt valid observations.
Evaluation emphasizes generalization across three dimensions. We construct a synthetic test set of 799 USGS sites (completely unseen during training) with expert-validated anomalies that are deliberately more complex than training corruptions, creating a four-axis separation (geographic, mathematical, temporal, parameter) that prevents pattern memorization. On this held-out set, HydroGEM achieves 0.70 F1 for anomaly detection and reduces reconstruction error by 70% relative to injected corruptions—though suggested corrections require expert review before operational use. For real-world validation, we evaluate zero-shot transfer to 100 Environment and Climate Change Canada sites using weak labels derived from operational corrections. Reconstructions show promising alignment with agency corrections in several cases, though direct comparison is limited by differences in correction methodology between USGS and ECCC workflows. This cross-national evaluation tests whether learned representations transfer beyond the training data source [kratzert2019towards, deznabi2024zero].
HydroGEM is designed for human-in-the-loop operational workflows [nearing2022lstm, goldstein2015peeking]. A three-tier flagging system passes high-confidence clean data with minimal review, prioritizes uncertain observations for expert inspection, and provides suggested reconstructions with uncertainty estimates for detected anomalies. Reconstructions are proposed corrections—not automated replacements—and require hydrologist approval before integration into official records. All outputs include provenance information suitable for audit and feedback. The goal is to expand hydrologists’ effective reach by automating routine screening and providing focused decision support, not to replace human expertise.
Contributions
This work advances hydrological quality control through:
  1. 1.
    A continental-scale foundation model trained on 3724 USGS sites with 6.67 million sequences—an order of magnitude larger than prior multi-site hydrological studies.
  2. 2.
    A two-stage training approach combining self-supervised pretraining on clean data with synthetic anomaly injection for detection and reconstruction, reducing dependence on labeled anomalies.
  3. 3.
    Hierarchical normalization enabling learning across six orders of magnitude while preserving scale-dependent physical behavior.
  4. 4.
    Expert-validated evaluation with 18 anomaly types and rigorous site-level partitioning ensuring zero-shot generalization assessment, achieving 0.70 F1 detection and 70% error reduction in suggested reconstructions.
  5. 5.
    Demonstrated cross-national transfer from USGS (USA) to ECCC (Canada) data, validating learned representations beyond training distribution.
Terminology Clean data: USGS approved records with quality code ‘A’ (Stage 1 training) Masked data: Clean data with random timesteps hidden for self-supervised reconstruction (Stage 1) Corrupted data: Clean data with synthetic anomalies injected (Stage 2 fine-tuning) Anomalous data: Real operational data requiring correction (deployment target) Suggested reconstruction: Model-proposed correction requiring hydrologist approval
2 Related Work
Our work bridges operational hydrological quality control, machine learning for time series anomaly detection, and foundation models in Earth system science. We review each area to position HydroGEM’s contributions.
2.1 Operational Hydrological Quality Control
Traditional quality control in hydrological monitoring combines manual expert review with rule-based automated checks [aquarius2024, schmidt2023saqc]. Commercial systems such as AQUARIUS (Aquatic Informatics) implement configurable range checks, rate-of-change thresholds, and statistical comparison filters, providing standardized workflows and audit trails [aquarius2024]. The open-source SaQC framework advances reproducibility through sequential test configurations and traceable quality flags [schmidt2023saqc].
However, rule-based approaches face fundamental limitations. Fixed thresholds require site-specific calibration and capture only obvious outliers. Context-dependent anomalies—such as ice effects that preserve plausible values but bias discharge [mcmillan2018uncertainty], gradual sensor drift within historical ranges [taormina2015neural], or rating curve shifts after channel-altering floods [addor2017large]—often evade detection. Multivariate relationships between discharge and stage, critical for identifying coupled sensor failures or unit conversion errors, are difficult to encode in rule-based systems. As monitoring networks expand and temporal resolution increases, manual calibration becomes prohibitively expensive.
Our approach addresses these limitations by learning multivariate patterns from continental-scale data, allowing context-aware detection without manual threshold tuning while preserving the human-in-the-loop workflows that agencies require.
2.2 Machine Learning for Time Series Anomaly Detection
Classical anomaly detection methods have been applied to environmental monitoring with mixed success. Statistical approaches (z-score tests, ARIMA residuals), distance-based methods (k-nearest neighbors, isolation forests), and one-class classifiers (one-class SVM, autoencoders) provide baseline capabilities [belay2023unsupervised, shi2023groundwater]. Recent environmental applications include wavelet-based multi-resolution analysis for river stage [nourani2014applications] and comparative evaluations of isolation forests versus one-class SVM for groundwater monitoring, where OCSVM achieved 88% precision on synthetic data [shi2023groundwater].
Deep learning approaches show stronger performance by capturing complex temporal dependencies. LSTM networks and Transformer attention mechanisms excel at modeling long-range patterns in time series [hochreiter1997long, he2019temporal]. Temporal convolutional networks enable efficient parallel processing with exponentially large receptive fields [han2023twawdlstm]. Autoencoders—including variational and LSTM-augmented variants—detect anomalies via reconstruction errors [osman2024variational, lee2024lstm]. Generative adversarial networks improve robustness through learned representations of normal data distributions [chen2025gasn]. Synthetic anomaly generation has emerged as a key technique for training when domain-specific labeled data is unavailable [kim2024enhancing].
Despite these advances, prior work suffers from two critical limitations for continental-scale deployment. First, models typically train and evaluate on individual sites or small regional datasets (often <50<50 sites), failing to capture the extreme hydrological heterogeneity spanning snowmelt-dominated montane systems to ephemeral desert washes to regulated lowland rivers. Transfer learning across such diverse regimes remains largely unexplored in hydrological anomaly detection. Second, most approaches require labeled training data—either manual annotations of anomalous periods or curated datasets separating sensor failures from physical phenomena. Such labels are scarce in operational hydrology, where agency quality flags reflect editorial decisions rather than ground truth [dogo2019survey, belay2023unsupervised].
Our two-stage approach resolves both limitations: self-supervised pretraining on massive unlabeled clean data learns general hydrological representations, while synthetic anomaly injection enables supervised fine-tuning without expensive manual labeling [dooley2023forecastpfn, kim2024enhancing].
2.3 Foundation Models in Earth System Science
Foundation models—large networks pretrained on massive datasets with self-supervised objectives, then transferred to diverse downstream tasks—have transformed Earth system science [nguyen2023climax, pathak2022fourcastnet, lam2023graphcast, bi2023accurate]. ClimaX demonstrated that transformers pretrained on heterogeneous climate reanalysis data with masked token prediction effectively transfer to forecasting, downscaling, and multi-variable prediction across spatiotemporal scales [nguyen2023climax]. Weather foundation models (FourCastNet [pathak2022fourcastnet], GraphCast [lam2023graphcast], Pangu-Weather [bi2023accurate]) now achieve medium-range forecast skill competitive with operational numerical weather prediction at orders of magnitude lower computational cost. Industrial-scale efforts such as IBM/NASA’s Prithvi WxC (2.3B parameters) [prithvi2024wxc] and Oak Ridge’s ORBIT (>100>100B parameters) [wang2024orbit] are extending this paradigm to multi-variable Earth system prediction.
In hydrology, Kratzert and colleagues pioneered multi-basin learning by training LSTM networks across hundreds of catchments using the CAMELS dataset [kratzert2019towards], showing that models learn generalizable runoff generation processes that transfer to ungauged basins. Subsequent work confirmed that diverse training across basins improves predictions despite regional hydrological variation [gauch2021proper]. Remote sensing foundation models pretrained on satellite imagery have shown strong transfer to environmental monitoring tasks [ali2025environmental]. Recent perspectives advocate for foundation models as assistive tools in hydrometeorology when developed with appropriate domain constraints [ali2025environmental].
However, existing foundation models in Earth science focus primarily on gridded reanalysis data (weather/climate) or basin-scale streamflow prediction. To our knowledge, no prior work has applied foundation model principles to operational quality control for in situ sensor networks. This gap reflects unique challenges:
  • Extreme scale heterogeneity: Six orders of magnitude in discharge within a single dataset, requiring specialized normalization approaches
  • Irregular multi-site time series: Non-gridded data with site-specific characteristics (rating curves, instrument types, operational protocols)
  • Real-time deployment constraints: Inference must operate without access to ground truth, using only observable quantities
  • Deploy-safe requirements: The model must not corrupt valid data—a failure mode unacceptable in operational settings
  • Human oversight mandates: Agencies require interpretable outputs, uncertainty quantification, and audit trails for regulatory compliance
2.4 Positioning of HydroGEM
HydroGEM integrates advances from multiple domains while addressing the distinct challenges of operational hydrological quality control. From foundation models [nguyen2023climax, pathak2022fourcastnet], we adopt self-supervised pretraining on massive unlabeled datasets. From large-sample hydrology [kratzert2019towards, gauch2021proper], we use multi-site learning for cross-basin transfer. From recent deep learning work [dooley2023forecastpfn, kim2024enhancing], we use synthetic anomaly injection for task-specific fine-tuning. From operational QC systems [aquarius2024, schmidt2023saqc], we incorporate human-in-the-loop workflows and audit requirements.
The result is a system that: (1) learns from 3,505 diverse sites without site-specific tuning, (2) handles six orders of magnitude through hierarchical normalization, (3) detects diverse anomaly types via synthetic training, (4) shows rigorous zero-shot generalization through site-level evaluation, (5) provides suggested reconstructions for hydrologist review, and (6) operates deploy-safe with human oversight. While individual components draw inspiration from prior work, their integration in a validated operational system for continental-scale hydrological quality control represents a new contribution bridging machine learning and environmental monitoring.
Table 1 contrasts HydroGEM with related approaches.
Table 1: Positioning HydroGEM relative to related work
Approach Sites Self-sup. Zero-shot Synthetic Recon. Deploy-safe
Rule-based QC [aquarius2024, schmidt2023saqc] Any N/A N/A No Manual Yes
Single-site ML [he2019temporal, lee2024lstm] 1–10 Varies No Sometimes Varies No
Multi-basin hydro [kratzert2019towards] 100–500 No Yes No No N/A
Weather FM [nguyen2023climax, pathak2022fourcastnet] Gridded Yes Yes No No N/A
HydroGEM 5,421 Yes Yes Yes Suggested Yes
3 Data and Evaluation Framework
Our evaluation strategy combines three data sources (Figure 1): (1) a continental-scale USGS corpus for foundation model pretraining, (2) synthetic test sets with expert-validated anomalies for controlled generalization assessment, and (3) real-world Canadian stations for zero-shot cross-national transfer validation. This multi-tiered design allows rigorous evaluation under both controlled conditions with known ground truth and operational settings with genuine quality control challenges.
Figure 1: Geographic distribution of study sites. USGS stations are split into training (n=3,724), validation (n=798), and test (n=799) sets with no geographic overlap. Canadian ECCC stations (n=100) were selected based on data completeness and quality criteria described in Section 3.4.2.
3.1 Foundation Training Corpus
3.1.1 USGS Streamflow Archive
We assembled a continental-scale dataset from 5,321 USGS streamgaging stations via the National Water Information System (NWIS) [usgs2024nwis], covering January 2000 through December 2024. For each station, we selected the best 10-year continuous period with high-quality paired discharge ($Q$) and stage ($H$) observations. All records carry USGS qualification code ‘A’ (approved), distinguishing them from provisional data subject to revision. Instantaneous readings (typically 15-minute intervals) were aggregated to hourly resolution via arithmetic means for consistent temporal scale.
The 10-year site-specific selection strategy balances data quality, temporal depth, and infrastructure consistency. Post-2000 focus ensures modern instrumentation across the network, while 10-year spans capture multiple seasonal cycles and extreme hydroclimatic events (droughts, floods) essential for robust anomaly detection. This period also corresponds to relatively standardized USGS quality control protocols, reducing confounding institutional factors.
3.1.2 Spatial and Hydrologic Coverage
The network spans the conterminous United States, Alaska, and Hawaii (162.69°W to 67.32°W, 19.71°N to 61.71°N), providing true continental-scale geographic diversity. Drainage areas range from 0.03 to 932,800 km2 (median 196 km2), consistent with scale-invariant river network organization. Gauge elevations span 69-69 m to 2,747 m (mean 467 m), capturing coastal, lowland, montane, and alpine monitoring sites.
This spatial coverage ensures exposure to principal North American flow generation mechanisms: snowmelt-dominated montane regimes, rainfall-dominated humid systems, mixed snow-rain transitional climates, ephemeral arid flows, and regulated downstream conditions [chow1988applied]. Such hydrologic diversity is critical for developing transferable representations that generalize beyond specific flow regimes.
3.1.3 Quality Control and Preprocessing
[…]
Gap filling uses hierarchical strategies: linear interpolation for gaps $\leq 6$ hours, exponential recession models [vogel1992baseflow, tallaksen1995hydrograph] for 6-24 hour gaps, and exclusion for gaps >24>24 hours. This protocol retains 94.7% of potential sequences, with 89.3% containing no interpolation. Windows intersecting flagged spans are dropped, eliminating information leakage.
3.1.4 Dataset Partitioning
To assess generalization, we assign entire sites to a single partition, avoiding leakage of site-specific signatures [kratzert2019towards]. The split allocates 70% training (3724 sites), 15% validation (798 sites), and 15% test (799 sites), preserving geographic coverage and proportional hydrologic region representation.
Continuous hourly series are segmented into 576-hour windows (24 days), capturing storm hydrographs, weekly patterns, and diurnal fluctuations while remaining tractable for transformers [vaswani2017attention]. Asymmetric stride—48 hours for training (dense overlap maximizing data utilization), 192 hours for validation/test (reducing autocorrelation)—produces 6.03M training sequences, 0.32M validation, and 0.33M test sequences (Table 2).
Table 2: Foundation corpus statistics
Partition Sites Windows (M) Length (h) Stride (h) Timesteps (M)
Training 3,724 6.03 576 48 3,470.59
Validation 798 0.32 576 192 185.98
Test 799 0.33 576 192 187.35
Total 5,321 6.67 576 3,843.92
3.2 Preprocessing Pipeline
3.2.1 Feature Engineering
At each timestep $t$ for site $s$, the model observes $\mathbf{x}_{t}\in\mathbb{R}^{12}$ combining static descriptors, dynamic hydrology, and derived temporal context (Table 3). This compact design balances information richness, computational tractability, and operational constraints—specifically, we exclude meteorological forcings (precipitation, temperature) to maximize applicability to gauge-only deployment scenarios.
Table 3: 12-dimensional feature set composition
Category Features Dimension
Static basin Latitude, longitude, drainage area, elevation 4
Dynamic hydrology Discharge ($Q$), stage ($H$) 2
Scale embeddings $\sigma_{\ln Q}$, $\sigma_{\ln H}$ (training std. dev. of log-transformed series) 2
Cross-site ranks Ordinal position: rank($A_{d}$)/NN, rank($z_{g}$)/NN 2
Seasonal context Monthly anomalies: $(Q-\mu_{Q,m})/\sigma_{Q,m}$, $(H-\mu_{H,m})/\sigma_{H,m}$ 2
Total 12
Static descriptors $\{\phi,\lambda,A_{d},z_{g}\}$ provide coarse hydrograph controls through latitude (climatic gradients), longitude (monsoon influence), drainage area (flashiness), and elevation (temperature lapse rates). Dynamic variables $\{Q,H\}$ constitute the core physical state, with stage encoding local geometry and backwater effects [nittrouer2013backwater] invisible in discharge alone. Derived features include scale embeddings returning absolute variability information lost during standardization, cross-site ranks stabilizing learning across magnitude orders, and monthly anomalies expressing seasonal departures.
3.2.2 Hierarchical Normalization
Hydrological magnitudes vary by six orders across sites (0.1 to 100,000 ft3/s), creating severe optimization challenges. Standard approaches fail: global standardization yields large-river dominance; site-specific alone loses cross-site comparability; min-max scaling suffers outlier sensitivity; raw units prevent convergence.
We introduce a three-tier hierarchical normalization ( full mathematics in Appendix A) simultaneously achieving: (1) stable gradients across extreme heterogeneity, (2) zero train-test leakage, (3) exact physical unit recovery, and (4) scale-dependent information preservation.
Tier 1 - Logarithmic stabilization: Apply $Q^{(1)}=\ln(Q+\epsilon)$, $H^{(1)}=\ln(H+\epsilon)$ for approximately log-normal variables [sangal1970lognormal], linearizing rating curves and stabilizing variance.
Tier 2 - Site-specific standardization: Compute $\mu_{s}$, $\sigma_{s}$ exclusively from each site’s training data, then transform $\mathbf{x}^{(2)}=(\mathbf{x}^{(1)}-\mu_{s})/(\sigma_{s}+\epsilon)$. For validation/test sites, substitute global training statistics, maintaining strict partition separation.
Tier 3 - Global clipping: Apply $\mathbf{x}^{\text{norm}}=\text{clip}(\mathbf{x}^{(2)},-3,+3)$ to inputs only, retaining 99.7% of normal variation while preventing gradient explosion. Outputs denormalize without clipping.
Exact inverse: $\hat{y}=\exp[\hat{y}^{\text{norm}}\cdot(\sigma_{s}+\epsilon)+\mu_{s}]-\epsilon$ recovers physical units. Scale embeddings ($\sigma_{\ln Q}$, $\sigma_{\ln H}$) return absolute variability information enabling scale-dependent behavior (e.g., flashiness in small basins).
Rationale: The logarithm converts multiplicative noise to additive; site-specific standardization enables weight sharing without large-river dominance; clipping provides numerical stability; embeddings distinguish 10 ft3/s (small creek) from 10,000 ft3/s (large river) despite identical post-normalization values. No single-tier normalization achieves all requirements for continental-scale learning.
3.3 Training-Time Anomaly Injection
Anomaly detection requires labeled corruption patterns, but manual annotation of 6M sequences is prohibitively expensive. We implement on-the-fly synthetic injection generating diverse patterns during stage 2 fine tuning for anomaly detection which is also a part of our training process.
Deliberate simplification philosophy: The training injector implements $\sim$11 simplified patterns (spikes, drift, flatlines, dropouts, saturation, clock shifts, quantization, unit jumps, warping, splicing) applied in normalized log-space, deliberately less sophisticated than test anomalies. This design forces learning of fundamental hydrometric consistency principles (discharge-stage coupling, temporal smoothness, physical plausibility) rather than memorizing specific signatures. If training and test used identical patterns, strong performance could arise from pattern matching rather than genuine understanding. The training-test complexity gap creates a defendable generalization test.
Controlled coverage: Two-tier system assigns light corruption (60% probability, 5-15% coverage) or moderate (40% probability, 15-30% coverage) with iterative refinement ($\pm 3\%$ tolerance for light, $\pm 5\%$ for moderate), maintaining mean 15.2% $\pm$ 3.1% across batches. Mixing: single-type (60%) teaches type-specific patterns; double-type (40%) teaches discrimination. Segments: $n\in[2,4]$ with lengths $L\in[T/100,T/4]$ randomly positioned. Curriculum scheduling ramps injection probability from 0.2 (epochs 1-2) to 0.4 (thereafter). Complete implementation details in Appendix E.
3.4 Evaluation Datasets
3.4.1 Synthetic Test Set Design
Rigorous generalization assessment requires test data preventing memorization while probing genuine abstraction. We construct a synthetic test set from 799 held-out USGS sites with expert-validated anomalies spanning 18 consolidated types (Table 5). Four orthogonal separation axes between training and test distributions ensure strong performance requires learning fundamental hydrometric principles rather than pattern matching.
Table 4: Four-axis training-test separation strategy
Axis Training Test Purpose
Geographic 3,724 sites 799 non-overlapping sites Eliminate site memorization
Mathematical Linear drift, basic offsets Exponential/sigmoid/polynomial; 3-4 variants per type Require functional abstraction
Temporal 8-96 hour segments Micro (3-58h), Meso (7-192h), Macro (72-520h) Probe scale-invariant detection
Parameter Coverage 10-32%, nominal strength Coverage trimodal: 3-9% / 32-44% / 44-60%; Severity bimodal Test distribution boundaries
Geographic separation: Zero site overlap eliminates memorization of station-specific patterns. Mathematical separation: Each test anomaly type implements 3-4 equation variants (e.g., drift: linear, exponential, sigmoid, polynomial) producing similar visual patterns through different mechanisms. Temporal separation: Structured duration regimes (micro/meso/macro) spanning 3-520 hours, with macro extended to 520h based on expert feedback regarding ice persistence [santeford1986ice]. Parameter separation: Coverage and severity anti-overlap the training sweet spot, testing distribution boundaries.
Expert validation and taxonomy consolidation.
We assembled an HTML validation gallery with 8+ representative sequences per anomaly type, reviewed by domain experts using structured protocols (realism scoring, temporal validation, confound detection). Reviewers included a geoscience faculty member with extensive experience in water resources research and a research assistant from the Water Resources Institute at the University of Vermont with operational USGS internship experience examining real-world data quality issues. Based on systematic feedback, we consolidated to 18 types with field-relevance agreement (Table 5), categorized by confidence: high (9 types: dropout, flatline, spike, backwater, ice_backwater, drift, rating_drift, debris_effect, sedimentation), medium (4 types), low (4 types), and debated (retained with caveat). Complete consolidation mapping and expert feedback integration details appear in Appendix G.
Table 5: Consolidated anomaly taxonomy (18 types)
Category Types Confidence Variants
Sensor failures Dropout, flatline, spike High 3-4 each
Hydraulic phenomena Backwater, ice_backwater, debris_effect, sedimentation High 3-4 each
Gradual degradation Drift, rating_drift, sensor_fouling High/Medium 3-4 each
Processing errors Bias_step, time_misalignment, quantization Medium/Low 3-4 each
Complex artifacts Splice, warp, noise_burst, gate_operation Low/Debated 3-4 each
Physical-space injection.
All test anomalies inject in denormalized physical space to respect actual discharge-stage coupling. The three-step process: (1) denormalize to physical units, (2) apply parametric transformations with hydraulic constraints (e.g., ice: $H^{\prime}=H(1+\alpha_{\text{ice}})$, $Q^{\prime}=Q(1-\beta_{\text{ice}})$ with $\alpha\sim U(0.15,0.55)$, $\beta\sim U(0,0.10)$), (3) renormalize with clipping. Single-type isolation (30% of sequences) enables per-type diagnostics; compound anomalies (70%, with 40% overlap probability) test discrimination. Complete injection implementation details, equation variants, and reproducibility protocols appear in Appendix F.
Rationale for synthetic evaluation.
Synthetic data provides defendable evaluation: deliberate separation prevents memorization; physical grounding ensures realism; multi-axis design probes distinct generalization modes; expert validation provides external validity; controlled conditions enable fine-grained diagnostics. Complementarity with operational validation (Section 3.4.2) provides stronger evidence than either alone—synthetic proves learned concepts, operational proves practical transfer.
3.4.2 Canadian Zero-Shot Transfer Evaluation
The USGS archive provides quality-controlled products but lacks systematic raw-vs-corrected record pairs exposing operational correction workflows. To test real-world transfer, we evaluate HydroGEM on Environment and Climate Change Canada (ECCC) stations distributing hourly raw and corrected unit values for both discharge and stage.
Zero-shot protocol: HydroGEM was trained exclusively on USGS sites and never fine-tuned on Canadian data. All experiments are strictly zero-shot, testing whether learned representations generalize across political boundaries, agencies, instrumentation protocols, and rating curve derivation methods.
[…]
Quality filtering: To ensure meaningful evaluation, we applied hydrologically-motivated checks to corrected records: sufficient variability (CV >0.10>0.10), monotonic rating ($\rho>0.5$), reasonable exponent ($b\in[0.5,10]$), moderate fit ($R^{2}\geq 0.3$), valid data fraction ($\geq 70\%$), limited flatlines ($30\%$ differences <0.001<0.001 ft). Station-level filtering excluded sites losing $5\%$ timesteps. Detailed filtering protocols appear in Appendix H.
Weak label construction: We compute relative changes induced by operational correction: $\Delta Q_{\text{rel}}=|Q_{\text{corr}}-Q_{\text{raw}}|/(|Q_{\text{raw}}|+\epsilon)$, similarly for $H$. Timesteps with $\Delta Q_{\text{rel}}>0.01$ or $\Delta H_{\text{rel}}>0.01$ are marked anomalous ($y_{\text{corr}}=1$), providing weak labels encoding where human expertise deemed raw data untrustworthy. The 1% threshold is intentionally conservative, capturing meaningful hydrograph adjustments.
These labels are “weak” because human editors may miss subtle anomalies or correct for operational reasons orthogonal to data quality (e.g., re-rating after channel changes). Nevertheless, they provide externally-validated ground truth unavailable in synthetic datasets. Pattern-based labels (local correlation deviations) complement correction-based labels, emphasizing multi-point segments where corrected hydrograph shape deviates from raw.
Site sampling: From all stations satisfying quality criteria, we randomly sampled 100 distinct stations for evaluation, ensuring representative out-of-sample behavior without cherry-picking. Window filtering required 10-40% correction fraction, identifying substantial but not overwhelming editing effort. For each site, we evaluate HydroGEM in pure zero-shot configuration, comparing binary anomaly masks against correction-based and pattern-based labels.
3.5 Summary
Our data preparation pipeline yields:
  • Foundation corpus: 6.67M sequences (3.84B hourly timesteps) from 5,008 USGS stations (70/15/15 train/val/test split)
  • Training anomalies: On-the-fly injection of $\sim$11 simplified types, 15.2% mean coverage, normalized-space corruption
  • Test anomalies: 18 expert-validated types, 4-axis separation, physical-space injection, 3-4 equation variants per type
  • Real-world transfer: 100 randomly sampled Canadian stations, weak labels from operational corrections, strict zero-shot evaluation
This multi-tiered strategy allows rigorous capability assessment: controlled synthetic tests probe specific generalization modes under known ground truth; real-world Canadian data validates operational transfer. Together, these provide evidence of learned hydrometric consistency principles rather than dataset artifacts.
4 HydroGEM Model Architecture
Having established the data preparation pipeline in Section 3, we now describe the HydroGEM model that processes the hierarchically normalized 12-dimensional feature vectors to detect and correct anomalies in streamflow time series (Figure 2). The model input $\mathbf{X}\in\mathbb{R}^{T\times 12}$ consists of the features defined in Section 3.2.1, transformed through the three-tier normalization scheme (Section 3.2.2), and segmented into 576-hour windows (Section 3.2.3). Each sequence represents a continental-scale sample spanning diverse hydrologic regimes, requiring the architecture to learn scale-invariant patterns while respecting physical constraints.
Figure 2: HydroGEM training and inference pipeline. Stage 1 pretrains a hybrid TCN-Transformer backbone (14.2M parameters) using masked reconstruction on clean USGS data. Stage 2 freezes the backbone and trains a lightweight detection head ( 11K parameters) on synthetically corrupted sequences. At inference, the model provides anomaly probabilities with uncertainty estimates via MC Dropout, and suggested corrections for human review.
4.1 Two-Stage Training Framework
HydroGEM uses a two-stage training approach that decouples representation learning from task-specific optimization, following recent advances in foundation model development [devlin2019bert, He2022]. This architectural design addresses the challenge of learning robust hydrological representations from limited labeled anomaly data while maintaining generalization across diverse monitoring sites and flow regimes.
The first stage trains a deep autoencoder backbone through self-supervised learning on clean hydrological sequences, allowing the model to discover core relationships between discharge, stage, basin characteristics, and temporal dynamics without requiring explicit anomaly labels. This pretraining phase processes the 6.03 million training sequences described in Section 3.2.3 to develop general-purpose hydrological representations that capture the full spectrum of normal flow behaviors across continental scales. The second stage augments this pretrained backbone with a specialized detection head and fine-tunes the system using the synthetically generated anomalies described in Section 3.3.2—approximately 11 simplified corruption patterns applied in normalized space with controlled coverage (mean 15.2% $\pm$ 3.1%).
This decoupled approach provides several advantages over end-to-end anomaly detection architectures. The self-supervised pretraining exploits massive volumes of unlabeled data that would be prohibitively expensive to manually annotate, learning subtle hydrological patterns that might not be captured from limited labeled examples. The pretrained representations show superior transfer learning capabilities [TransferLearning2021], allowing detection of new anomaly types not explicitly included in the fine-tuning dataset. Furthermore, the modular design permits independent optimization of reconstruction quality and detection performance, facilitating ablation studies and architectural improvements without complete retraining.
4.2 Stage 1: Self-Supervised Pretraining
4.2.1 Backbone Network Architecture
The backbone combines Temporal Convolutional Networks (TCNs) with Transformer-based attention in a hierarchical encoder-decoder structure. This hybrid approach exploits complementary strengths: TCNs for efficient multi-scale local pattern extraction and Transformers for long-range dependencies. Complete architectural equations appear in Appendix B.
TCN Encoder.
The encoder maps 12-dimensional inputs to 128-dimensional hidden representations through four stacked TCN blocks with exponentially increasing dilation rates ($r\in\{1,2,4,8\}$). Each block uses residual connections, batch normalization, and dropout ($p=0.2$). With kernel size $k=3$, this yields receptive fields of 5, 13, 29, and 61 hours for blocks 1-4, capturing hydrological processes from sub-daily fluctuations to multi-day storm events. The TCN encoder contributes 2.1 million parameters.
Transformer.
The TCN output is projected from 128 to 256 dimensions and processed through 4 transformer layers with 8-head attention. We use Cosine Retention Attention with learnable temporal decay [Mongaras2024] to prevent attention collapse in long sequences. Sliding window attention (256 positions) reduces complexity from $O(T^{2})$ to $O(T\cdot W)$. The transformer contributes 9.8 million parameters.
Decoder and Skip Connections.
The decoder mirrors the encoder, with four TCN blocks reconstructing the original sequence. A gated skip connection adaptively combines encoder and decoder pathways, creating a curriculum learning effect: early training relies on skip connections for gradient flow while later training strengthens the transformer pathway. The gate parameter $\alpha$ is initialized to yield near-zero contribution from the skip path, gradually increasing during training.
The complete backbone comprises 14.2 million parameters: TCN encoder (2.1M), transformer (9.8M), TCN decoder (2.1M), and projection layers (0.2M).
4.2.2 Masking Strategy and Objectives
The masking strategy in Stage 1 serves a fundamentally different purpose than the anomaly injection in Stage 2 (Section 3.3). While Stage 2 trains the model to detect corrupted values, Stage 1 establishes a robust representation of normal hydrological behavior through reconstruction of missing data. This pretraining creates a strong inductive bias—the model learns the statistical regularities, physical constraints, and temporal dynamics that characterize clean hydrological data. When subsequently exposed to anomalies during Stage 2, the model can detect them precisely because they violate these learned patterns of normality.
The distinction between missing and corrupted data is critical. Missing data (Stage 1) requires the model to leverage contextual information and physical relationships to infer likely values, teaching it the underlying structure of hydrological systems. Corrupted data (Stage 2) contains incorrect values that must be identified and corrected, a task made possible by the model’s pretrained understanding of what constitutes normal behavior. This two-stage approach has proven successful in vision [He2022] and language understanding [devlin2019bert], and we adapt it here for hydrological quality control.
We implement four masking patterns calibrated to the typical data gaps encountered in operational monitoring, informed by masking strategies in self-supervised learning [Wettig2023]:
  • Point masking (40% probability): Randomly masks 15% of timesteps, forcing reconstruction from immediate temporal context. This teaches the model short-term hydrological persistence and smoothness constraints.
  • Block masking (30% probability): Masks 1-3 contiguous blocks of 12-72 hours, requiring the model to learn recession curves, diurnal patterns, and multi-day correlations essential for bridging extended gaps.
  • Periodic masking (20% probability): Masks 4-hour windows at 168-hour intervals, teaching the model to recognize weekly operational cycles and anthropogenic patterns in streamflow.
  • Feature masking (10% probability): Masks either discharge (70%) or stage (30%) for 24-168 hours, forcing the model to learn rating curve relationships [Kennedy1984] and cross-variable dependencies.
Each training sequence undergoes masking with probability 0.80. The resulting reconstruction task requires the model to internalize three critical aspects of hydrological data: (1) temporal dynamics including recession rates and hydrograph shapes, (2) physical relationships between discharge and stage governed by hydraulic geometry, and (3) scale-dependent patterns that vary with basin size and flow regime. These learned representations provide the foundation for detecting anomalies that violate expected patterns during operational deployment.
4.2.3 Pretraining Loss Functions
The pretraining objective combines five loss components (complete formulations in Appendix C):
  • Weighted reconstruction loss: Prioritizes discharge (weight 3.0) and stage (2.5) over other features
  • Temporal consistency loss: Preserves recession slopes and rising limb characteristics
  • Variance preservation loss: Prevents over-smoothing to collapsed mean predictions
  • Scale consistency loss: Maintains denormalization capability for physical unit recovery
  • Diversity regularization: Prevents attention collapse through entropy maximization
The backbone trains for 20 epochs using AdamW optimization [Loshchilov2017] with OneCycleLR scheduling (peak learning rate $5\times 10^{-4}$), gradient clipping (norm 1.0), and early stopping (patience 7 epochs). Training requires approximately 48 GPU-hours on NVIDIA A100 hardware.
4.3 Stage 2: Anomaly Detection Fine-Tuning
The second stage augments the pretrained backbone with a specialized detection head and fine-tunes the complete system using the synthetic anomaly injection framework described in Section 3.3. As detailed in that section, the training injector implements approximately 11 simplified anomaly patterns (spikes, drift, flatlines, dropouts, saturation, clock shifts, quantization, unit jumps, warping, splicing, and subtle drift) applied in normalized log-space with controlled coverage targeting 15.2% $\pm$ 3.1% temporal coverage. This deliberate simplification—using simple corruption mechanisms during training while testing on complex physical-space anomalies (Section 3.4)—forces the model to learn fundamental hydrometric consistency principles rather than memorizing specific anomaly signatures.
4.3.1 Detection Head Architecture
The detection head operates entirely on observable quantities without requiring ground-truth clean data during inference [Angiulli2023, Gong2019]. It computes an 11-dimensional feature vector from the relationship between potentially corrupted observations and backbone reconstructions:
  • Reconstruction residuals: Absolute differences between observed and reconstructed discharge/stage
  • Temporal gradients: Forward differences revealing drift and trend anomalies
  • Rolling variability: 7-hour window statistics distinguishing sensor malfunction from natural variability
  • Rating curve deviation: Local power-law fit quality indicating hydraulic inconsistency
  • Cross-correlation features: Coupled anomalies affecting both variables simultaneously
Features undergo robust standardization using median absolute deviation and are processed through a two-layer MLP (128 and 64 units) with GELU activation and dropout ($p=0.2$). Monte Carlo dropout (10 forward passes) provides uncertainty quantification for operational decision-making.
4.3.2 Fine-Tuning Objectives
The fine-tuning loss balances four objectives (complete formulations in Appendix C):
  • Focal loss: Addresses class imbalance with $\alpha=0.25$, $\gamma=2.0$
  • Corruption reconstruction: Ensures accurate corrections on anomalous segments
  • Clean preservation: Penalizes modifications to uncorrupted data (essential for operational trust)
  • Physics constraints: Enforces discharge-stage coupling and rating curve compliance
Fine-tuning proceeds for 12 epochs with AdamW ($lr=5\times 10^{-5}$), requiring approximately 8 GPU-hours.
4.4 Inference Pipeline
At deployment, the trained model processes incoming observations through the frozen backbone to generate reconstructions, then computes detection features from observation-reconstruction residuals (Figure 2, bottom). The detection head outputs anomaly probabilities with uncertainty estimates obtained via MC Dropout ($N=20$ forward passes). High-uncertainty predictions are flagged for mandatory expert review. For detected anomalies, the backbone reconstruction serves as a suggested correction—not an automated replacement—requiring hydrologist approval before integration into official records. This human-in-the-loop design ensures data quality while enabling the model to screen thousands of stations continuously.
5 Experimental Setup
5.1 Computational Infrastructure
All experiments were run on the Lonestar6 supercomputer at the Texas Advanced Computing Center, accessed through the NSF National AI Research Resource (NAIRR) Pilot program. The backbone pretraining phase ran for 20 epochs and required approximately 46 hours on NVIDIA A100 GPUs. The fine tuning phase completed in 12 epochs. For both stages we selected the checkpoint with the lowest validation loss and used that model state for all downstream evaluation.
5.2 Evaluation Metrics
We evaluate HydroGEM along 2 complementary dimensions: anomaly detection performance and reconstruction quality.
For anomaly detection we report precision, recall, F1 score, and the area under the receiver operating characteristic curve (AUROC). Precision measures the fraction of flagged points that are true anomalies, recall measures the fraction of true anomalies that are correctly flagged, and F1 is the harmonic mean of precision and recall. F1 is our primary reporting metric, with AUROC providing a threshold independent view of ranking quality.
For reconstruction quality we focus on settings where a clean reference is available. On the synthetic USGS test set we compute the relative reduction in absolute error between raw anomalous values and model reconstructions,
$\text{Error Reduction}=\frac{\lvert X^{(\text{raw})}-X^{(\text{clean})}\rvert-\lvert\hat{X}-X^{(\text{clean})}\rvert}{\lvert X^{(\text{raw})}-X^{(\text{clean})}\rvert}\times 100\%,$ (1)
and report segment wise averages together with root mean squared error (RMSE) on anomalous segments. We additionally track RMSE on non anomalous timesteps to confirm that the model does not degrade clean observations. On the synthetic experiments HydroGEM typically reduces error by around 70% on corrupted segments, while keeping errors on clean segments close to the baseline.
For the Canadian case study, reconstructions are evaluated more qualitatively. Weak labels are derived from operational corrections, and correction strategies, especially for ice affected periods where discharge is often pushed very low while stage remains relatively high, differ from the USGS corrections used to train HydroGEM. Official guidance notes that stage affected by backwater from ice may accurately represent local conditions but may not be suitable for discharge computation, and operational products explicitly warn that ice can raise stage for a given flow [usgs_wdr_docs, eccc_data_manual, eccc_ice_warning]. In our sample many Environment and Climate Change Canada winter corrections reduce discharge much more aggressively than the USGS style corrections seen in training, particularly during ice control periods where stage remains elevated. As a result, in the Canadian evaluation we treat reconstruction error as a secondary diagnostic rather than a primary performance metric and place emphasis on detection metrics.
5.3 Baseline Methods
To assess zero shot anomaly detection performance we compare HydroGEM against 12 baseline methods that do not require labeled training data or site specific calibration. All baselines operate in a zero shot mode at new stations, matching HydroGEM’s deployment setting. We consider 3 categories.
Statistical baselines.
These methods represent simple distributional checks that are widely used in practice. They include a Z Score detector that flags observations with absolute z score greater than 3 for either discharge or stage, an Interquartile Range (IQR) rule that flags values outside the interval $Q_{1}-1.5\times\text{IQR},Q_{3}+1.5\times\text{IQR}]$, and a Moving Average Residual detector that uses a centered 7 day window and flags points whose residuals exceed 3 standard deviations of the rolling window. Together these capture point wise outliers and deviations from local temporal context.
Unsupervised machine learning.
We include generic unsupervised anomaly detectors that require no labels and can be applied at scale. Isolation Forest and Local Outlier Factor (LOF) operate on standardized discharge and stage and use standard contamination settings, representing tree based and density based approaches to unsupervised outlier detection. A Seasonal Trend decomposition using LOESS (STL) baseline models a 7 day seasonal component and flags residuals that exceed 3 standard deviations. These methods provide a representative sample of modern off the shelf anomaly detectors that do not use hydrological structure.
Hydrological domain methods.
The third group encodes relationships that hydrologists routinely use during manual quality control. A Rating Curve Residual baseline fits a power law relationship $Q=a(H-H_{0})^{b}$ by log-linear regression and flags points with large log-space residuals. This serves as a key reference because rating residuals are central to operational practice for detecting rating shifts, drift, and ice effects. Additional checks include rate-of-change limits on discharge and stage, persistence detection for stuck sensors, Q-H consistency checks that flag periods where discharge and stage trends diverge or correlations break down, and a seasonal envelope that flags values outside month-specific percentile bands. Together these methods approximate the mix of physical rules and heuristics that agencies use in production systems.
Methods not included.
We intentionally exclude methods that rely on extensive labeled anomalies or site specific tuning. Supervised deep learning classifiers and autoencoders require labels for each station and would not represent a true zero shot setting. Per site reconstruction models that are fit directly on test windows are also excluded, since they would not be deployed in this way operationally. Proprietary automated flagging systems from agencies such as USGS and Environment and Climate Change Canada are not reproducible without internal configuration parameters. Pretraining multiple alternative deep architectures on the full USGS corpus would be methodologically interesting but is beyond the present computational scope and is left for future work.
5.4 Evaluation Protocols
Synthetic USGS evaluation.
For synthetic evaluation we use a held out set of 799 USGS sites that are completely unseen during training. Clean sequences from these sites are corrupted with injected anomalies that cover a broad range of types, durations, spatial coverage, and severities. We report aggregate detection metrics over all anomalies and also stratify performance by anomaly type and basic regime attributes such as duration and coverage. Detailed breakdowns are provided in the appendix, while the main text focuses on overall detection performance and the relative ranking of HydroGEM against baselines.
Reconstruction quality on this synthetic set is summarized by the error reduction and RMSE metrics described above, computed only on timesteps where injected anomalies are present, with separate reporting of RMSE on clean timesteps.
Canadian zero shot evaluation.
For cross national evaluation we apply HydroGEM without retraining to 100 Canadian stations from Environment and Climate Change Canada. Weak labels are derived from human corrections in the operational record. These labels reflect editorial decisions and local correction practice, so we report F1 and related detection metrics over a range of labeling thresholds defined by the relative magnitude of human edits, with full curves in the appendix. This provides a sensitivity view of how performance changes as the notion of meaningful correction becomes stricter. Given the documented differences between USGS and Environment and Climate Change Canada correction strategies in ice affected periods [usgs_wdr_docs, eccc_data_manual, eccc_ice_warning], we interpret reconstruction errors here as supportive evidence rather than a primary outcome and focus our conclusions on detection performance.
6 Results
We evaluate HydroGEM through two complementary assessments: (1) detection and reconstruction on a synthetic test set with exact ground-truth labels, and (2) zero-shot transfer to Canadian stations where labels derive from operational hydrologist corrections. The synthetic evaluation establishes baseline capabilities under controlled conditions, while the Canadian evaluation tests generalization to real-world data quality challenges across national boundaries.
6.1 Synthetic Anomaly Detection
6.1.1 Overall Detection Performance
Table 6 and Figure 3 present detection performance for HydroGEM and eleven baseline methods on the synthetic test set. HydroGEM achieves F1 = 0.792 (precision: 0.755, recall: 0.832), substantially outperforming all baselines. The strongest baseline, Isolation Forest, attains F1 = 0.392, representing a 101.9% relative improvement by HydroGEM. All pairwise improvements are statistically significant ($p<0.001$, Wilcoxon signed-rank test).
Table 6: Detection performance on the synthetic test set. All baseline thresholds were derived from site-specific data statistics following QARTOD methodology, ensuring no parameter tuning on test data.
Method Category F1 Precision Recall
HydroGEM (Ours) Foundation Model 0.792 0.755 0.832
Isolation Forest Unsupervised ML 0.392 0.600 0.291
IQR Statistical 0.146 0.621 0.082
Q-H Consistency Hydrological 0.118 0.608 0.065
STL Residual Unsupervised ML 0.117 0.747 0.064
LOF Unsupervised ML 0.100 0.563 0.055
Seasonal Envelope Hydrological 0.078 0.649 0.041
Rating Curve Hydrological 0.056 0.703 0.029
Z-Score Statistical 0.037 0.665 0.019
Rate of Change Hydrological 0.032 0.767 0.016
Moving Average Statistical 0.015 0.645 0.008
Persistence Hydrological 0.001 0.004 0.001
Figure 3: Overall detection performance on the synthetic test set. HydroGEM (F1 = 0.792) outperforms all baselines, with Isolation Forest (F1 = 0.392) as the nearest competitor.
The baseline hierarchy reveals distinct performance tiers. Isolation Forest (F1 = 0.392) achieves the strongest baseline performance by leveraging multivariate structure in the discharge-stage feature space. Traditional statistical methods (Z-Score, IQR, Moving Average) achieve F1 << 0.15, as they operate on univariate signals and cannot capture stage-discharge coupling violations. Domain-specific approaches like Q-H Consistency (F1 = 0.118) and Rating Curve residuals (F1 = 0.056) underperform despite their hydrological grounding—these methods detect only anomalies that violate instantaneous rating relationships, missing temporally-extended distortions that preserve local Q-H ratios.
6.1.2 Per-Anomaly-Type Analysis
Figure 4 presents F1 scores stratified by anomaly type and detection method. HydroGEM achieves the highest F1 for all 18 anomaly types, demonstrating consistent generalization across the full spectrum of hydrometric failures.
Figure 4: Detection F1 scores by anomaly type and method. HydroGEM achieves the highest performance across all 18 anomaly types. Isolation Forest provides the strongest baseline overall, particularly for dropout and unit mismatch detection.
Performance varies systematically with anomaly characteristics. HydroGEM achieves strongest detection for backwater effects (F1 = 0.77), ice backwater (F1 = 0.76), and drift anomalies (F1 = 0.73)—conditions that produce sustained Q-H decoupling patterns. Moderate performance occurs for sensor-level artifacts including bias steps (F1 = 0.61), flatlines (F1 = 0.63), and spikes (F1 = 0.61), which manifest as localized discontinuities rather than extended temporal patterns. The most challenging anomaly type is rating shift (F1 = 0.51), where abrupt but physically-plausible changes in the stage-discharge relationship can resemble legitimate rating curve updates.
The heatmap also reveals baseline specializations. Isolation Forest achieves its best performance on dropout (F1 = 0.44) and unit mismatch (F1 = 0.46), both of which produce outliers in multivariate feature space. However, no baseline achieves F1 >> 0.20 for subtle anomalies like sensor fouling, splice artifacts, or time misalignment—patterns that require learning complex temporal dependencies rather than applying fixed decision rules.
6.1.3 Reconstruction Quality
Beyond detection, HydroGEM provides corrected time series through its reconstruction head. On the synthetic test set, the model achieves 68.7% mean error reduction, computed as:
$\text{Error Reduction}=1-\frac{\text{MAE}(\hat{x},x_{\text{clean}})}{\text{MAE}(x_{\text{corrupt}},x_{\text{clean}})}$ (2)
where $\hat{x}$ denotes the HydroGEM reconstruction, $x_{\text{clean}}$ the ground-truth clean signal, and $x_{\text{corrupt}}$ the corrupted input.
Reconstruction performance correlates with anomaly severity and duration. For high-magnitude anomalies (>>25% deviation from clean values), error reduction reaches 74.2%. Medium-duration events (6–48 hours) achieve 71.3% reduction, compared to 64.1% for short events (<<6 hours) where limited temporal context constrains inference. The model preserves clean data with high fidelity: on uncorrupted segments, reconstruction MAE remains below 2% of the signal range, confirming that corrections are applied selectively.
6.1.4 Representative Detection Examples
Figure 5 illustrates successful detection on a challenging multi-anomaly window containing overlapping backwater effects, exponential drift, and debris-induced artifacts spanning approximately 400 hours. HydroGEM achieves F1 = 0.810 (precision: 0.995, recall: 0.683) with 63.7% error reduction. The high precision indicates minimal false alarms on clean segments, while the reconstruction closely tracks the ground-truth signal through complex, overlapping corruption patterns.
Figure 5: Successful detection example with overlapping backwater, exponential drift, and debris effects. HydroGEM achieves F1 = 0.810 with 63.7% error reduction. Top panels show stage and discharge time series (green: clean, red dashed: corrupted, orange: HydroGEM corrected). Bottom panel compares ground-truth labels with model detections.
Figure 6 presents a challenging case containing multiple gate operation events producing rapid, physically-valid discharge fluctuations. HydroGEM achieves perfect recall (1.000) but lower precision (0.622), resulting in F1 = 0.767 with slightly negative error reduction (-3.2%). The model flags legitimate gate-induced transients as suspicious—these rapid Q-H excursions resemble sensor artifacts in the learned representation. This case illustrates the inherent difficulty of distinguishing anthropogenic flow modifications from sensor malfunctions without explicit metadata about control structures.
Figure 6: Challenging case with gate operation events. HydroGEM achieves F1 = 0.767 with perfect recall but reduced precision, as legitimate gate-induced transients produce Q-H patterns similar to sensor artifacts. This motivates treating model outputs as suggestions requiring expert review.
This failure mode motivates an important operational consideration: HydroGEM outputs should be treated as quality control suggestions rather than autonomous corrections, particularly for stations with known flow regulation. The model’s high recall ensures that genuine anomalies are rarely missed, while expert review can filter false positives arising from unusual but valid hydrological conditions.
6.2 Zero-Shot Transfer to Canadian Stations
The ultimate test of a foundation model is generalization beyond its training distribution. We evaluated HydroGEM on 100 stations from Environment and Climate Change Canada’s (ECCC) hydrometric network—data the model never encountered during training. This assessment tests whether representations learned from USGS stations capture universal principles of streamflow monitoring that transfer across national boundaries, instrumentation practices, and climatic regimes.
6.2.1 Baseline Comparison
Table 7 presents detection performance on Canadian stations. We applied the same eleven baseline methods using identical QARTOD-derived threshold protocols described in Section 6.1.1. HydroGEM achieves F1 = 0.586 at the 1% correction threshold, substantially outperforming all baselines with improvements statistically significant at $p<0.001$ (Wilcoxon signed-rank test).
Table 7: Detection performance on Canadian zero-shot evaluation (F1 at 1% correction threshold).
Method Category F1 Precision Recall
HydroGEM Foundation Model 0.586 0.544 0.635
Persistence Hydrological 0.418 0.428 0.410
Isolation Forest Unsupervised ML 0.263 0.354 0.210
IQR Statistical 0.104 0.139 0.086
LOF Unsupervised ML 0.101 0.096 0.107
STL Residual Unsupervised ML 0.058 0.111 0.039
Seasonal Envelope Hydrological 0.049 0.082 0.035
Rate of Change Hydrological 0.038 0.067 0.027
Rating Curve Hydrological 0.028 0.036 0.034
Z-Score Statistical 0.028 0.031 0.032
Q-H Consistency Hydrological 0.024 0.022 0.027
Moving Average Statistical 0.009 0.012 0.007
Among baselines, persistence detection performs best (F1 = 0.418), reflecting that stuck sensors constitute a meaningful fraction of operational corrections in the Canadian archive. However, persistence detection is narrowly specialized: it identifies only zero-variance intervals and cannot detect rating curve violations, ice effects, sensor drift, or backwater conditions. Isolation Forest ranks second (F1 = 0.263), consistent with its general-purpose outlier detection capability. Statistical methods and most hydrological heuristics achieve F1 << 0.11, indicating that simple threshold-based rules inadequately capture the complexity of real-world data quality issues.
The baseline ranking differs notably between synthetic and Canadian evaluations. Isolation Forest performs strongly on synthetic data (F1 = 0.392) but less so on Canadian data (F1 = 0.263), while Persistence shows the opposite pattern (synthetic: F1 = 0.001; Canadian: F1 = 0.418). This divergence reflects differences in anomaly composition: synthetic anomalies include diverse corruption patterns where multivariate outlier detection excels, whereas Canadian operational corrections predominantly address sensor stalls and ice effects that Persistence can partially identify through its stuck-sensor detection. HydroGEM’s consistent strong performance across both evaluation settings shows robustness to these distributional differences.
6.2.2 Representative Detection Example
Figure 7 illustrates HydroGEM’s detection behavior on station 05AD028 during a spring period encompassing ice-affected and open-water conditions. The station achieved F1 = 0.859, demonstrating strong agreement with operational corrections. During clean periods (late March), the model correctly preserves data integrity with minimal false positives. The major stage spike on April 2 is appropriately flagged and reconstructed.
Figure 7: Detection example for Canadian station 05AD028 (F1 = 0.859). Blue and red shading indicate periods where ground truth and HydroGEM identify anomalies, respectively. The model preserves clean observations (late March) while flagging data quality issues. Reconstruction divergence during ice-affected periods (early March) reflects differences between USGS training corrections and ECCC operational practices [Turcotte2013, Beltaos2013].
The ice-affected period (March 16–24, blue shading) exhibits reconstruction divergence where ECCC operators applied aggressive discharge reductions while stage remained elevated—a characteristic ice-backwater correction strategy. HydroGEM’s proposed reconstructions reflect discharge-stage relationships learned from USGS training data, which differ systematically from Canadian practices in ice-control periods. This divergence does not indicate detection failure; rather, it shows that the model learned USGS correction principles rather than memorizing site-specific patterns. For Canadian deployment, we recommend treating reconstructions as suggestions requiring operator review, particularly during ice-affected periods.
6.2.3 Detection Behavior Characterization
Beyond aggregate metrics, we analyzed whether HydroGEM learned physically meaningful patterns that transfer to Canadian conditions. Figure 8 summarizes detection behavior across seasons, correction magnitudes, and anomaly durations.
Seasonal alignment.
We compared HydroGEM’s flag rate against the human correction rate across seasons (Figure 8A). Both exhibit consistent seasonal structure: winter shows the highest activity (HydroGEM: 53.6%, human: 68.2%), followed by spring (50.6%, 59.5%), fall (42.9%, 56.4%), and summer (39.3%, 42.4%). The winter peak reflects well-documented challenges in cold-region hydrometry: ice-affected flow produces backwater conditions violating standard rating curves, anchor ice and frazil ice cause sensor interference, and instrument freezing creates data gaps [WMO2010a, Turcotte2013, Beltaos2013]. That HydroGEM—trained on predominantly ice-free USGS stations—identifies these signatures in Canadian data suggests the model learned generalizable representations of sensor malfunction and rating curve violation.
Magnitude-independent detection.
A concern with learned detectors is whether they simply flag large deviations while missing subtle corrections. Table 8 stratifies detection recall by correction magnitude. HydroGEM maintains consistent recall (0.48–0.56) across corrections ranging from minor adjustments (1–5%) to major revisions (50–100%), indicating the model learned anomaly patterns—temporal dynamics, channel inconsistencies, physical constraint violations—rather than relying on deviation magnitude as a proxy.
Table 8: Detection recall by correction magnitude. Consistent performance across magnitude ranges indicates pattern-based rather than magnitude-based detection.
Correction Magnitude Timesteps Detected Recall
1–5% 209,223 114,903 0.52
5–10% 54,174 32,297 0.49
10–25% 69,868 36,054 0.48
25–50% 63,022 33,030 0.52
50–100% 141,958 77,670 0.56
Temporal scale coverage.
Operational data quality issues span diverse temporal scales. Analysis of contiguous anomaly segments reveals this heterogeneity: 48% of correction events span fewer than 6 hours (sensor spikes, transient failures), 29% span 6–48 hours (storm-event artifacts, equipment malfunctions), and 23% exceed 48 hours (extended ice cover, prolonged sensor failures). HydroGEM’s training on synthetic anomalies spanning multiple temporal scales allows detection across this full spectrum (Figure 8C).
Figure 8: Detection behavior on Canadian zero-shot evaluation. (A) Seasonal comparison of HydroGEM flag rate versus human correction rate, showing aligned structure with winter peaks. (B) Per-site agreement between model and human corrections. (C) Distribution of anomaly event durations. (D) Detection recall by correction magnitude.
6.3 Summary of Results
Across both evaluation settings, HydroGEM shows strong anomaly detection performance with substantial margins over established baselines. On synthetic data with exact ground truth, the model achieves F1 = 0.792 with 68.7% reconstruction error reduction, outperforming the best baseline by 36.3%. On Canadian stations in zero-shot transfer, HydroGEM achieves F1 = 0.586, exceeding the best baseline by 40.2%. The model generalizes across all 18 expert-validated anomaly types, aligns with operational correction patterns across seasons, and maintains sensitivity to both subtle and severe anomalies.
The architectural separation between training anomalies (simplified patterns in normalized space) and test anomalies (physically-grounded injection with multiple equation variants) confirms that these results reflect learned hydrometric principles rather than pattern memorization. Zero-shot transfer to Canadian stations—with different instrumentation, correction practices, and ice-dominated hydrology—provides additional evidence of generalizable representations. Complete per-window results and visualizations are available at https://github.com/[REPO]/hydrogem-results.
6.4 Ablation Studies
We conducted systematic ablation studies to validate key design decisions. These experiments, performed on validation subsets during development, reveal the necessity of each architectural component for achieving strong zero-shot generalization.
6.4.1 Normalization Strategy Impact
The hierarchical normalization scheme emerged from systematic failures when applying standard approaches to continental-scale data. Figure 9 shows convergence behavior across different normalization strategies on a 100-site validation subset.
Figure 9: Training dynamics under different normalization strategies. (a) Convergence comparison showing only hierarchical normalization achieves stable optimization. (b) Performance across stream size categories reveals severe bias in global standardization toward large rivers.
Raw units cause gradient explosion as discharge values spanning six orders of magnitude create unstable optimization. Global standardization converges but exhibits severe bias: small streams achieve F1 = 0.31 while large rivers reach F1 = 0.69, indicating the model essentially ignores small-scale hydrology. Site-specific normalization alone prevents cross-site learning, achieving mediocre uniform performance (F1 $\approx$ 0.52) across all scales. Min-max normalization fails to converge as single outlier events repeatedly reshape the normalization range.
Our hierarchical approach (log transform $\rightarrow$ site standardization $\rightarrow$ global clipping with scale embeddings) represents the minimal configuration achieving three requirements simultaneously: stable gradients across training, meaningful cross-site comparison, and preservation of absolute scale information through embeddings. The scale embeddings ($\sigma_{\ln Q}$, $\sigma_{\ln H}$) prove particularly critical, reintroducing magnitude information that enables scale-appropriate anomaly detection thresholds.
6.4.2 Architectural Component Analysis
The hybrid TCN-Transformer architecture addresses fundamental limitations of using either component in isolation. Figure 10 illustrates performance degradation at different temporal scales when architectural components are removed.
Figure 10: Architectural ablation results. (a) Detection performance by anomaly duration shows TCN-only degradation beyond its 61-hour receptive field, while Transformer-only fails on short events. (b) Training curves demonstrate 2x slower convergence without gated skip connections.
The TCN-only variant maintains strong performance (F1 >> 0.70) for anomalies under 24 hours but degrades precipitously for extended events. With receptive field mathematically limited to 61 hours (4 blocks, maximum dilation 8), the model cannot capture multi-week sensor drift or seasonal rating changes. Performance drops below F1 = 0.40 for anomalies exceeding 336 hours.
Conversely, the Transformer-only variant shows opposite behavior: poor performance on short transients (F1 = 0.45 for 1-6 hour events) but improving with duration. Self-attention struggles to isolate brief spikes within 576-hour windows, effectively smoothing high-frequency patterns that constitute many operational anomalies.
The gated skip connection implements crucial curriculum learning. Without it, convergence requires approximately 20 epochs versus 10 with the skip path, and final reconstruction error remains 35% higher. The learnable gate parameter evolves from near-zero initially ($\sigma(\alpha)\approx 0.007$) to moderate values ($\approx 0.35$) by training completion, progressively incorporating learned representations while maintaining gradient flow.
6.4.3 Training Strategy Validation
Table 9 quantifies the benefit of two-stage training through controlled experiments on 50-site validation subsets. Each variant used identical architecture and hyperparameters, differing only in training methodology.
Table 9: Training strategy comparison on validation subset. Metrics averaged across three random seeds with standard deviation.
Training Strategy Detection F1 Clean Preservation Convergence
Two-stage (ours) 0.684 ± 0.021 96.8% 12 epochs
End-to-end synthetic 0.521 ± 0.034 89.3% 8 epochs
No pretraining 0.398 ± 0.048 91.2% 15+ epochs
Pretrain only (no finetune) 0.216 ± 0.019 98.1% N/A
End-to-end training with synthetic anomalies from random initialization shows 31% lower F1 and higher variance, suggesting overfitting to specific corruption patterns. The model learns to detect training anomalies but fails to generalize to novel corruption types. Training without pretraining yields poorest detection performance and highest variance, as the model lacks robust priors about normal hydrological behavior.
The pretrained backbone provides stable initialization with learned representations of recession curves, rating relationships, and seasonal patterns. This foundation enables rapid fine-tuning convergence and superior final performance while maintaining 96.8% clean data preservation—critical for operational trust.
6.4.4 Loss Component Importance
Each pretraining loss component prevents specific failure modes. Figure 11 shows validation error when individual components are removed, ordered by impact severity.
Figure 11: Loss component ablation ordered by impact. Removing temporal consistency causes 42% degradation, while scale consistency loss increases error by 33%. Each component addresses operational requirements.
Temporal consistency loss proves most critical, with removal increasing MAE by 42%. Without this constraint, reconstructions exhibit physically implausible discontinuities—sudden jumps in discharge while stage remains constant—that would trigger numerous false positives during detection. Scale consistency loss (33% degradation) enables accurate denormalization to physical units, essential for providing meaningful corrections to operators. Variance preservation prevents over-smoothed predictions that would miss legitimate flow variability.
The weighted reconstruction loss focusing on discharge (weight = 3.0) and stage (weight = 2.5) over static features reflects operational priorities. Uniform weighting degrades performance by 21%, particularly for discharge reconstruction where accuracy is paramount for water resource management decisions.
6.4.5 Synthesis of Ablation Findings
These ablations demonstrate that HydroGEM’s performance emerges from synergistic interaction between components rather than any single innovation:
  • Hierarchical normalization allows learning across six orders of magnitude while preserving scale-aware detection capabilities
  • Hybrid architecture captures complementary temporal scales, with TCN handling short transients and Transformer modeling extended dependencies
  • Two-stage training prevents overfitting to synthetic patterns while establishing robust hydrological priors
  • Multi-component loss ensures physically plausible reconstructions meeting operational requirements
Performance degrades 20-42% when removing major components, confirming the current architecture represents a minimal configuration for continental-scale deployment. Each design choice addresses specific failure modes encountered when scaling anomaly detection beyond individual sites to thousands of heterogeneous monitoring stations.
7 Discussion
7.1 Interpretation of Key Findings
The successful development and validation of HydroGEM shows that foundation model principles can be effectively adapted to operational hydrological quality control, addressing a critical bottleneck in environmental monitoring infrastructure. The results reveal several important insights about both the model’s capabilities and the nature of streamflow data quality challenges.
The two-stage training approach proves remarkably effective at learning generalizable hydrometric principles without extensive labeled anomaly datasets. By pretraining on 6.03 million clean sequences, HydroGEM develops robust representations of normal hydrological behavior that serve as an inductive bias for anomaly detection. Recession curves, rating relationships, and seasonal patterns emerge naturally from the self-supervised objective. The subsequent fine-tuning with simplified synthetic anomalies forces the model to learn fundamental consistency principles rather than memorizing specific corruption signatures. The four-axis separation between training and test distributions provides strong evidence that the 36.3% performance gain over baselines reflects genuine abstraction rather than pattern matching. This finding suggests that the historical challenge of limited labeled anomaly data in operational hydrology can be circumvented through careful architectural design and training strategies.
The hierarchical normalization scheme successfully allows learning across six orders of magnitude in discharge, from ephemeral desert washes to major rivers, within a single model. This addresses a fundamental challenge in continental-scale hydrology where traditional approaches either require site-specific calibration or suffer from large-river dominance. The combination of logarithmic stabilization, site-specific standardization, and scale embeddings preserves both local patterns and absolute magnitude information. The model can thus distinguish a 10 ft3/s anomaly in a small creek from a 10,000 ft3/s anomaly in a major river, despite their identical appearance in normalized space. This capability is essential for continental-scale deployment where manual calibration of thousands of sites would be prohibitively expensive.
Zero-shot transfer to Canadian stations shows that learned representations capture universal principles of streamflow monitoring rather than USGS-specific artifacts. The alignment between HydroGEM’s detection patterns and operational correction rates across seasons is particularly striking. The winter peak in both model flagging and human corrections, corresponding to documented ice effects [Turcotte2013, Beltaos2013], emerges despite training predominantly on ice-free stations. This suggests the model learned to identify fundamental signatures of sensor malfunction and rating curve violation that manifest similarly across different hydrological contexts. The ability to generalize across national boundaries, instrumentation protocols, and climatic regimes validates the foundation model approach for operational hydrology.
7.2 Limitations and Operational Considerations
Despite strong performance, several limitations require careful consideration for operational deployment. These constraints arise from both technical architecture decisions and fundamental challenges in defining “correct” hydrological data.
Reconstruction divergence in ice-affected periods represents the most significant operational challenge. While HydroGEM successfully detects ice-related anomalies with high recall, reconstructions during these periods often diverge from operational corrections, particularly for Canadian stations. This divergence reflects fundamental differences in correction philosophy between agencies. USGS guidance emphasizes maintaining physical plausibility of both discharge and stage, typically applying moderate adjustments that preserve the coupled relationship. In contrast, ECCC practice often aggressively reduces discharge to near-zero while preserving elevated stage readings during ice backwater, acknowledging that stage may accurately represent local conditions even when unsuitable for discharge computation [eccc_ice_warning]. This is not a detection failure but rather highlights that “correct” values depend on operational context, downstream use cases, and agency-specific protocols. For deployment, we strongly recommend treating all reconstructions as suggestions requiring expert review, particularly during known ice periods.
The model occasionally flags rapid but valid discharge fluctuations from gate operations, hydropeaking, or flash floods as anomalous. These events produce sudden excursions in discharge-stage relationships that resemble sensor artifacts in the learned representation. While maintaining high recall ensures genuine anomalies are rarely missed (a critical requirement for operational systems), this behavior reinforces the necessity of human oversight. The challenge of distinguishing anthropogenic flow modifications from sensor failures without explicit metadata represents a fundamental limitation of approaches based solely on time series patterns. Future work could incorporate auxiliary information about control structures, precipitation events, or operational schedules to improve discrimination.
Training requires substantial computational resources, with backbone pretraining consuming approximately 48 GPU-hours on modern hardware. While this one-time cost amortizes across thousands of deployment sites, it may limit accessibility for smaller agencies or research groups. Inference remains efficient at less than 100ms per window on standard hardware, enabling real-time deployment. However, the current architecture processes 576-hour windows in batch mode, creating a tradeoff between temporal context and streaming responsiveness. Operational systems requiring immediate anomaly detection would benefit from architectural modifications enabling incremental updates while maintaining long-term memory.
The model operates solely on observable quantities without meteorological forcing, maximizing applicability to gauge-only stations but preventing discrimination between sensor failures and physical extremes using precipitation context. A 500% discharge spike could represent sensor malfunction or an actual flash flood; without rainfall data, the model cannot definitively distinguish these cases. Integration with gridded meteorological products or real-time weather station networks could improve discrimination, though at the cost of increased system complexity and reduced deployability to remote sites lacking complementary observations.
7.3 Advancing the Field
HydroGEM advances hydrological quality control beyond previous approaches along multiple dimensions. Traditional rule-based systems require extensive manual calibration and capture only simple threshold violations. Even the strongest baseline overall, Isolation Forest, achieves only F1 = 0.392 on synthetic data despite being a general-purpose unsupervised anomaly detector. Moreover, baseline performance varies dramatically between evaluation settings: Isolation Forest drops to F1 = 0.263 on Canadian data, while Persistence rises from F1 = 0.001 (synthetic) to F1 = 0.418 (Canadian), highlighting how specialized methods fail when anomaly distributions shift. Rating curve residual methods, despite their theoretical grounding in hydraulic principles, achieve F1 = 0.056 on synthetic data because they examine only instantaneous relationships while missing temporal evolution patterns. HydroGEM maintains strong performance across both evaluation settings through learned representations that adapt to local context while capturing both instantaneous and temporal dependencies.
Prior machine learning approaches in hydrology have typically focused on individual sites or small regional clusters, requiring site-specific training that limits scalability. Studies reporting strong performance often evaluate on single watersheds or validate using random train-test splits that allow information leakage between partitions. By training on 3,724 sites and evaluating on 799 completely unseen locations with rigorous geographic separation, we establish that multi-site learning with appropriate normalization achieves better generalization than specialized models. The zero-shot transfer to Canadian stations provides additional evidence that continental-scale training produces robust representations applicable beyond the training distribution.
Within the broader foundation model framework, HydroGEM represents the first systematic application to in situ environmental sensor networks. Weather and climate foundation models have achieved remarkable success with gridded reanalysis data, but sensor networks present unique challenges including irregular spatial coverage, extreme magnitude heterogeneity, sparse labeling, and mandatory human oversight requirements. Our two-stage training framework and hierarchical normalization scheme provide a template for adapting foundation model principles to these operational constraints. The approach could generalize to other environmental monitoring applications including water quality sensors, air quality networks, and ecological observation systems.
7.4 Broader Implications for Operational Hydrology
The development of HydroGEM has implications extending beyond immediate quality control applications. As monitoring networks expand globally, driven by decreasing sensor costs and climate adaptation needs, the gap between data generation and quality assurance capacity will continue widening. Current estimates suggest the global water monitoring network generates over 100 million observations daily, yet most water agencies employ fewer than 50 hydrologists for quality control. HydroGEM offers a path to scale expert oversight by automating routine screening while preserving human judgment for complex decisions. The three-tier flagging system enables efficient triage, passing high-confidence clean data with minimal review, prioritizing uncertain observations for inspection, and providing suggested corrections for detected anomalies.
The human-AI collaboration approach embodied by HydroGEM may serve as a model for AI integration in safety-critical environmental monitoring. Rather than pursuing full automation, the system amplifies human expertise through scalable pattern recognition while preserving accountability, interpretability, and corrective capacity. All outputs include provenance information suitable for regulatory audit. This design philosophy acknowledges that data quality decisions often involve value judgments about intended use, acceptable uncertainty, and operational constraints that extend beyond pure statistical pattern recognition.
Climate change intensifies the importance of reliable hydrological observations for adaptation planning, infrastructure design, and ecosystem management. As extreme events become more frequent and hydrological regimes shift beyond historical baselines, distinguishing sensor anomalies from genuine extremes becomes increasingly critical. HydroGEM’s ability to detect subtle violations of learned patterns while adapting to local context could help identify both data quality issues and emerging hydrological signatures of environmental change. The pretrained backbone could be fine-tuned for specific climate impact detection tasks, potentially serving as an early warning system for regime shifts.
8 Conclusion
This work introduced HydroGEM, a foundation model for continental-scale streamflow quality control that addresses the growing mismatch between automated data generation and expert review capacity in hydrological monitoring networks. Through careful architectural design and training strategies, we showed that foundation model principles can be successfully adapted to operational sensor networks despite challenges including extreme heterogeneity, sparse labeling, and stringent deployment constraints.
The technical contributions of this work center on three innovations. First, the two-stage training approach combines self-supervised pretraining on massive unlabeled data with synthetic anomaly fine-tuning, eliminating dependence on scarce labeled datasets while learning robust representations of hydrological normality. Second, the hierarchical normalization scheme allows learning across six orders of magnitude within a single model, addressing a key challenge in continental-scale hydrology. Third, the rigorous evaluation framework with expert-validated synthetic anomalies and zero-shot cross-national transfer provides strong evidence for genuine generalization rather than pattern memorization.
Experimental validation confirms the effectiveness of these approaches. On synthetic test data with known ground truth, HydroGEM achieves F1 = 0.792 with 68.7% reconstruction error reduction, outperforming the strongest baseline by 36.3% while maintaining 97% preservation of clean data. Zero-shot transfer to 100 Canadian stations yields F1 = 0.586, with detection patterns aligning closely with operational correction rates across seasons. The model successfully identifies all 18 expert-validated anomaly types and maintains consistent performance across correction magnitudes from 1% to 100%, indicating pattern-based rather than threshold-based detection.
Several avenues for future development emerge from this work. Immediate priorities include incorporating meteorological forcing to provide physical context for extreme values, developing streaming architectures for real-time deployment with incremental updates, and creating specialized fine-tuning procedures for ice-dominated watersheds where correction practices vary significantly between agencies. The modular architecture enables these extensions without complete retraining. Longer-term opportunities include multi-modal fusion with satellite observations for comprehensive watershed monitoring, extension to water quality parameters where sensor drift and biofouling present similar challenges, and development of active learning frameworks that continuously improve from operational corrections.
As environmental monitoring transitions from sparse manual sampling to dense automated sensing, quality control emerges as the critical bottleneck preventing full utilization of these data streams for scientific understanding and operational decision-making. HydroGEM shows that foundation models can help bridge this gap through human-AI collaboration that preserves expert oversight while enabling continental-scale monitoring. This approach, where AI handles pattern recognition at scale while humans provide domain expertise and accountability, may prove essential as we build the observational infrastructure needed to understand and manage water resources under accelerating environmental change.
Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. EAR 2012123. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Any use of trade, firm, or product names is for descriptive purposes only and does not imply endorsement by the U.S. Government.
The work was also supported by the University of Vermont College of Engineering and Mathematical Sciences through the REU program.
We acknowledge the US Geological Survey (USGS) and Environment and Climate Change Canada (ECCC) for making the hydrometric datasets publicly available. We also thank David Baude, Research Analyst at the Water Resources Institute, University of Vermont, for his expert review and valuable feedback on the synthetic anomaly taxonomy.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability
USGS streamflow data are publicly available through the National Water Information System (NWIS) at https://waterdata.usgs.gov/nwis. Environment and Climate Change Canada hydrometric data are available at https://wateroffice.ec.gc.ca/. Processed datasets and code will be made available upon publication at [repository URL to be added].
CRediT Author Statement
Ijaz Ul Haq: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing - Original Draft, Visualization. Byung Suk Lee: Supervision, Resources, Writing - Review & Editing, Funding acquisition. Julia N. Perdrial: Supervision, Domain expertise, Writing - Review & Editing, Funding acquisition.
Appendix A Hierarchical Normalization
This appendix provides complete mathematical derivations and implementation details for the three-tier hierarchical normalization scheme introduced in Section 3.2.2.
A.0.1 Tier 1: Logarithmic Stabilization
For approximately log-normal hydrological variables [sangal1970lognormal], we apply:
$Q^{(1)}=\ln(Q+\epsilon),\quad H^{(1)}=\ln(H+\epsilon)$ (3)
with $\epsilon=10^{-8}$ to handle near-zero flows in ephemeral systems.
Rationale: This transformation achieves three critical objectives:
1. Rating curve linearization.
Natural channel rating curves follow power-law form $Q=a(H-H_{0})^{b}$ [chow1988applied]. In log-space:
$\ln Q\approx b\ln H+\ln a$ (4)
This linearization enables the network to learn discharge-stage coupling through simple linear relationships rather than complex nonlinear functions.
2. Multiplicative-to-additive noise conversion.
Hydrological measurement errors are often proportional to magnitude (e.g., 5% uncertainty applies to both 10 ft3/s and 10,000 ft3/s). In linear space, this multiplicative noise has magnitude-dependent variance:
$\text{Var}(Q_{\text{measured}})=(0.05\cdot Q_{\text{true}})^{2}$ (5)
The log transformation converts multiplicative noise to additive noise with constant variance:
$\text{Var}(\ln Q_{\text{measured}})\approx\text{const}$ (6)
3. Heteroscedasticity stabilization.
Without log transformation, gradient magnitudes for a 10% error on a large river exceed those for small streams by $10^{3}$-$10^{6}\times$, preventing convergence. The log transformation ensures equal-percentage errors produce equal-magnitude gradients regardless of absolute scale.
A.0.2 Tier 2: Site-Specific Standardization
For each site $s$, we compute normalization statistics exclusively from that site’s training partition:
$\mu_{s}=\mathbb{E}_{t\in\text{train}(s)}[\mathbf{x}_{t}^{(1)}],\quad\sigma_{s}=\text{Std}_{t\in\text{train}(s)}[\mathbf{x}_{t}^{(1)}]$ (7)
and transform:
$\mathbf{x}_{t}^{(2)}=\frac{\mathbf{x}_{t}^{(1)}-\mu_{s}}{\sigma_{s}+\epsilon}$ (8)
with $\epsilon=10^{-8}$ for numerical stability.
Critical implementation detail - preventing data leakage: For sites appearing only in validation or test partitions, we substitute global training statistics computed over all training sites:
$\mu_{\text{global}}=\mathbb{E}_{s\in\text{train}}\mathbb{E}_{t\in\text{train}(s)}[\mathbf{x}_{t}^{(1)}],\quad\sigma_{\text{global}}=\text{Std}_{s\in\text{train}}\text{Std}_{t\in\text{train}(s)}[\mathbf{x}_{t}^{(1)}]$ (9)
This ensures strict partition separation—validation and test data never influences normalization parameters. Most multi-site hydrological models violate this principle by computing statistics over all sites (including validation/test), creating subtle information leakage.
Rationale: Site-specific standardization enables weight sharing across basins by placing each site’s typical range in a common normalized space, preventing large rivers from dominating the loss landscape. A small stream with $Q\in[1,10]$ ft3/s and the Mississippi River with $Q\in[10^{4},10^{6}]$ ft3/s both normalize to $\approx[-2,+2]$, contributing equally to gradient updates.
A.0.3 Tier 3: Global Clipping
To downweight extreme leverage points (measurement spikes, brief sensor failures, telemetry errors), we clip standardized variables to a symmetric bound:
$\mathbf{x}_{t}^{\text{norm}}=\text{clip}(\mathbf{x}_{t}^{(2)},-\tau,+\tau),\quad\tau=3.0$ (10)
Critical asymmetry: Clipping applies only to inputs during training. Predicted outputs are denormalized without clipping to avoid artificially truncating physical predictions.
Rationale: For normally distributed standardized variables, $\tau=3.0$ retains approximately 99.7% of values while capping extreme outliers. In long 576-hour sequences, even rare extreme values can cause gradient explosion. Clipping provides numerical stability while preserving the network’s ability to reconstruct full-range physical magnitudes through the exact inverse mapping.
A.0.4 Exact Inverse Mapping
Let $\hat{y}^{\text{norm}}\in\mathbb{R}^{12}$ denote a model output on the normalized scale. Recovery to physical units follows the reverse transformation chain:
$\hat{y}^{(1)}$ $=\hat{y}^{\text{norm}}\cdot(\sigma_{s}+\epsilon)+\mu_{s}$ (11)
$\hat{y}$ $=\exp(\hat{y}^{(1)})-\epsilon$ (12)
where Eq. 11 reverses standardization and Eq. 12 reverses log transformation.
Special cases:
  • If site $s$ used global statistics during normalization (validation/test site), apply the same global statistics in the inverse
  • Static variables never log-transformed (latitude, longitude, elevation, drainage area) skip Eq. 12
Numerical verification: We verify exact invertibility by round-trip transformation:
$\|\mathbf{x}_{\text{original}}-\text{denorm}(\text{norm}(\mathbf{x}_{\text{original}}))\|_{\infty}<10^{-6}$ (13)
for all training, validation, and test sequences.
A.0.5 Scale Embeddings
The scale embeddings $\{\sigma_{\ln Q,s},\sigma_{\ln H,s}\}$ complete the normalization system by returning information about absolute variability lost during standardization:
$\sigma_{\ln Q,s}=\text{Std}_{t\in\text{train}(s)}[\ln(Q_{t}+\epsilon)],\quad\sigma_{\ln H,s}=\text{Std}_{t\in\text{train}(s)}[\ln(H_{t}+\epsilon)]$ (14)
These embeddings are broadcast across the temporal dimension and concatenated with the normalized features $\mathbf{x}^{\text{norm}}$, enabling the network to condition on absolute scale.
Rationale: After site-specific standardization, a normalized discharge value of +1.0+1.0 is ambiguous—it could represent 10 ft3/s at a small headwater stream or 10,000 ft3/s at a large river. Both appear identical in normalized space. The scale embeddings resolve this ambiguity, enabling the network to express scale-dependent behavior (e.g., flashiness in small basins, baseflow dominance in large rivers) despite operating in normalized space.
Design choice validation: Ablation studies (Section 6.4) show that removing scale embeddings degrades performance by 12-15% F1 score, confirming their necessity for continental-scale learning.
Appendix B Architecture Equations
This appendix provides complete mathematical formulations for the TCN-Transformer backbone described in Section 4.2.1.
B.1 TCN Encoder
The dilated convolution operation for sequence $\mathbf{x}$ with filter $\mathbf{f}$ and dilation rate $r$:
$(\mathbf{x}*_{r}\mathbf{f})(t)=\sum_{i=0}^{k-1}f(i)\cdot x_{t-r\cdot i}$ (15)
Linear projection from input to hidden dimension:
$\mathbf{H}^{(0)}=\mathbf{X}\mathbf{W}_{\text{proj}}+\mathbf{b}_{\text{proj}},\quad\mathbf{W}_{\text{proj}}\in\mathbb{R}^{12\times 128}$ (16)
TCN block with residual connection:
$\text{TCN-Block}(\mathbf{H},r)=\mathbf{H}+\mathcal{F}(\mathbf{H},r)$ (17)
where $\mathcal{F}(\mathbf{H},r)=\text{Conv1D}_{r}(\text{ReLU}(\text{BN}(\text{Conv1D}_{r}(\mathbf{H}))))$.
Receptive field at block $\ell$:
$\text{RF}_{\ell}=2^{\ell+2}-3$ (18)
B.2 Transformer
Projection to transformer dimension:
$\mathbf{Z}^{(0)}=\mathbf{H}^{(4)}\mathbf{W}_{\text{up}}+\mathbf{b}_{\text{up}},\quad\mathbf{W}_{\text{up}}\in\mathbb{R}^{128\times 256}$ (19)
Query, key, value projections for head $h$:
$\mathbf{Q}_{h}=\mathbf{Z}\mathbf{W}_{Q,h},\quad\mathbf{K}_{h}=\mathbf{Z}\mathbf{W}_{K,h},\quad\mathbf{V}_{h}=\mathbf{Z}\mathbf{W}_{V,h}$ (20)
L2 normalization for cosine similarity:
$\hat{\mathbf{q}}_{i}=\frac{\mathbf{q}_{i}}{\|\mathbf{q}_{i}\|_{2}+\epsilon},\quad\hat{\mathbf{k}}_{j}=\frac{\mathbf{k}_{j}}{\|\mathbf{k}_{j}\|_{2}+\epsilon}$ (21)
Attention scores with temporal decay and causal mask:
$\mathbf{A}_{ij}=\frac{\hat{\mathbf{q}}_{i}^{T}\hat{\mathbf{k}}_{j}}{\sqrt{d_{h}}}\cdot\gamma^{|i-j|}\cdot\mathbb{1}_{i\geq j}$ (22)
Multi-head attention:
$\text{MultiHead}(\mathbf{Z})=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{8})\mathbf{W}_{O}$ (23)
Feed-forward network:
$\text{FFN}(\mathbf{x})=\text{GELU}(\mathbf{x}\mathbf{W}_{1}+\mathbf{b}_{1})\mathbf{W}_{2}+\mathbf{b}_{2}$ (24)
Layer normalization:
$\text{LayerNorm}(\mathbf{x})=\frac{\mathbf{x}-\mu}{\sqrt{\sigma^{2}+\epsilon}}\cdot\gamma+\beta$ (25)
B.3 Decoder and Skip Connections
Decoder projection:
$\mathbf{G}^{(0)}=\mathbf{Z}^{(L)}\mathbf{W}_{\text{down}}+\mathbf{b}_{\text{down}},\quad\mathbf{W}_{\text{down}}\in\mathbb{R}^{256\times 128}$ (26)
Final output projection:
$\hat{\mathbf{X}}_{\text{decoder}}=\mathbf{G}^{(4)}\mathbf{W}_{\text{out}}+\mathbf{b}_{\text{out}},\quad\mathbf{W}_{\text{out}}\in\mathbb{R}^{128\times 12}$ (27)
Gated skip connection:
$\hat{\mathbf{X}}=(1-\sigma(\alpha))\cdot\hat{\mathbf{X}}_{\text{decoder}}+\sigma(\alpha)\cdot\text{Linear}(\mathbf{H}^{(4)}_{\text{encoder}})$ (28)
Appendix C Loss Functions
C.1 Pretraining Loss
Combined pretraining objective:
$\mathcal{L}_{\text{pretrain}}=\mathcal{L}_{\text{recon}}+0.6\mathcal{L}_{\text{temporal}}+0.4\mathcal{L}_{\text{variance}}+0.3\mathcal{L}_{\text{scale}}+0.05\mathcal{L}_{\text{diversity}}$ (29)
Weighted reconstruction loss:
$\mathcal{L}_{\text{recon}}=\frac{1}{T}\sum_{t=1}^{T}\sum_{f=1}^{d}w_{f}\cdot(\hat{x}_{t,f}-x_{t,f})^{2}$ (30)
with weights $w=3.0$ (discharge), $2.5$ (stage), $1.5$ (seasonal), $1.0$ (static).
Temporal consistency loss:
$\mathcal{L}_{\text{temporal}}=\frac{1}{T-1}\sum_{t=1}^{T-1}\sum_{f\in\{Q,H\}}w_{f}\cdot[(\hat{x}_{t+1,f}-\hat{x}_{t,f})-(x_{t+1,f}-x_{t,f})]^{2}$ (31)
Variance preservation loss:
$\mathcal{L}_{\text{variance}}=\sum_{f=1}^{d}[\text{Var}(\hat{\mathbf{x}}_{f})-\text{Var}(\mathbf{x}_{f})]^{2}$ (32)
Scale consistency loss:
$\mathcal{L}_{\text{scale}}=\frac{1}{T}\sum_{t=1}^{T}\sum_{f\in\text{scale}}(\hat{x}_{t,f}-x_{t,f})^{2}$ (33)
Diversity regularization:
$\mathcal{L}_{\text{diversity}}=-\text{Entropy}(\mathbf{A})+\lambda_{\text{rank}}\max(0,10-\text{rank}(\mathbf{H}))$ (34)
C.2 Fine-Tuning Loss
Focal loss for detection:
$\mathcal{L}_{\text{focal}}=-\frac{1}{T}\sum_{t=1}^{T}\alpha_{t}(1-\hat{p}_{t})^{\gamma}[y_{t}\log\hat{p}_{t}+(1-y_{t})\log(1-\hat{p}_{t})]$ (35)
with $\alpha=0.25$, $\gamma=2.0$.
Corruption reconstruction loss:
$\mathcal{L}_{\text{corrupt}}=\frac{1}{|\mathcal{T}_{\text{anom}}|}\sum_{t\in\mathcal{T}_{\text{anom}}}\|\hat{\mathbf{X}}_{t}-\mathbf{X}_{t}^{(\text{clean})}\|_{2}^{2}$ (36)
Clean preservation loss:
$\mathcal{L}_{\text{preserve}}=\frac{1}{|\mathcal{T}_{\text{clean}}|}\sum_{t\in\mathcal{T}_{\text{clean}}}\|\hat{\mathbf{X}}_{t}-\mathbf{X}_{t}^{(\text{obs})}\|_{2}^{2}$ (37)
Physics constraint loss:
$\mathcal{L}_{\text{physics}}=\frac{1}{T-1}\sum_{t=1}^{T-1}\text{ReLU}(-\nabla_{t}Q\cdot\nabla_{t}H)+\lambda_{\text{RC}}\mathbb{E}[d_{\text{RC}}]$ (38)
Combined fine-tuning objective:
$\mathcal{L}_{\text{finetune}}=1.5\mathcal{L}_{\text{focal}}+\lambda_{c}(\epsilon)\mathcal{L}_{\text{corrupt}}+\lambda_{p}(\epsilon)\mathcal{L}_{\text{preserve}}+0.1\mathcal{L}_{\text{physics}}$ (39)
with $\lambda_{c}(\epsilon)=0.25\cdot\min(1,\epsilon/3)$ and $\lambda_{p}(\epsilon)=0.05\cdot\min(1,\epsilon/3)$.
Appendix D Quality Control Protocols
This appendix details the multi-tier quality control protocol applied to ensure training data integrity (Section 3.1.3).
D.0.1 Outlier Detection
Values exceeding four standard deviations from site-specific monthly means are flagged for manual review:
$\text{flag}(Q_{t})=\mathbb{I}\left[|Q_{t}-\mu_{Q,m(t),s}|>4\sigma_{Q,m(t),s}\right]$ (40)
where $m(t)$ denotes the month of timestep $t$, and $\mu_{Q,m,s}$, $\sigma_{Q,m,s}$ are computed from all observations in month $m$ for site $s$ within the training period.
Rationale: The adaptive monthly threshold accounts for natural hydrologic variability (e.g., higher variance during spring snowmelt) while identifying extreme leverage points. Fixed global thresholds would flag normal high flows in flashy basins while missing subtle anomalies in stable systems.
D.0.2 Physical Plausibility Checks
Temporal consistency.
We test rate-of-change using a site-specific threshold:
$\frac{|Q_{t}-Q_{t-1}|}{Q_{t-1}}>\theta_{\text{site}}$ (41)
where $\theta_{\text{site}}$ is computed as the 99th percentile of observed fractional changes during the training period for that site. This adaptive threshold accommodates both flashy small basins and slowly-responding large systems.
Rating curve validation.
[…]
Range bounds.
All values are checked against physically plausible ranges:
  • Discharge: $Q\in[0,Q_{\max}]$ where $Q_{\max}=\max(\text{training data})\times 2.0$
  • Stage: $H\in[H_{\min}-1.0,H_{\max}+1.0]$ ft, allowing 1-foot exceedance beyond observed training range
D.0.3 Gap Filling Strategy
Short gaps (1-6 hours): Linear interpolation
$Q_{t}=Q_{t_{0}}+(Q_{t_{1}}-Q_{t_{0}})\cdot\frac{t-t_{0}}{t_{1}-t_{0}}$ (42)
where $t_{0}$ and $t_{1}$ are the last valid timestep before and first valid timestep after the gap.
Rationale: Linear interpolation is appropriate for slowly-varying baseflow conditions where discharge changes smoothly. For 1-6 hour gaps, this assumption typically holds.
Medium gaps (6-24 hours): Exponential recession model
For gaps exceeding 6 hours but less than 24 hours, we apply the exponential recession relationship [vogel1992baseflow, tallaksen1995hydrograph]:
$Q(t)=Q(t_{0})\exp[-k(t-t_{0})]$ (43)
The recession constant $k$ is estimated via:
$k=-\frac{1}{\Delta t}\ln\left(\frac{Q(t_{0}+\Delta t)}{Q(t_{0})}\right)$ (44)
using a 48-hour window before the gap. If insufficient pre-gap data exists, we use the site-specific median recession constant computed from all valid recession segments in the training period.
Rationale: The exponential form respects fundamental watershed storage dynamics: $\frac{dS}{dt}=-kS$, where storage $S$ is linearly related to discharge. This is more physically realistic than linear interpolation for medium-duration gaps where recession behavior dominates.
Long gaps (>>24 hours): Exclusion
Gaps exceeding 24 hours are excluded entirely from sequence construction. No imputation is attempted to prevent long-range artifacts that could introduce spurious patterns.
D.0.4 Gap Statistics
The complete gap-filling protocol retains 94.7% of potential 576-hour sequences:
  • 89.3% contain no interpolation
  • 9.8% contain 1-3 interpolated hours (linear)
  • 0.9% contain 4-6 hours (mix of linear and recession)
  • 0% contain >>6 hours (sequences intersecting gaps >>6h are dropped)
Leakage prevention: Any sequence window intersecting a flagged timestep (outlier, physical implausibility, or gap >>6h) is excluded entirely, even if only a single hour is affected. This conservative approach eliminates potential information leakage from imputation.
Appendix E Training-Time Anomaly Injection
This appendix provides implementation details for the on-the-fly synthetic anomaly injection during training (Section 3.3).
E.0.1 Simplified Anomaly Types
The training injector implements approximately 11 simplified corruption patterns applied in normalized log-space:
  1. 1.
    Spike: Add Gaussian noise impulses: $x^{\prime}_{t}=x_{t}+\mathcal{N}(0,\alpha\sigma_{x})$ at random positions, $\alpha\sim U(2,5)$
  2. 2.
    Drift: Apply linear trend: $x^{\prime}_{t}=x_{t}+\beta\cdot(t-t_{\text{start}})$ where $\beta\sim U(-0.01,+0.01)$
  3. 3.
    Flatline: Freeze values: $x^{\prime}_{t}=x_{t_{\text{freeze}}}$ for all $t$ in segment
  4. 4.
    Dropout: Replace with near-zero: $x^{\prime}_{t}=\epsilon$ where $\epsilon\sim U(10^{-6},10^{-4})$
  5. 5.
    Saturation: Clip to limits: $x^{\prime}_{t}=\text{clip}(x_{t},x_{\min}+0.1,x_{\max}-0.1)$
  6. 6.
    Clock shift: Temporal offset: $x^{\prime}_{t}=x_{t+\Delta t}$ where $\Delta t\sim\{-3,-2,-1,+1,+2,+3\}$ hours
  7. 7.
    Quantization: Discretize: $x^{\prime}_{t}=\text{round}(x_{t}/\Delta q)\cdot\Delta q$ where $\Delta q\sim U(0.05,0.2)$
  8. 8.
    Unit jump: Abrupt bias: $x^{\prime}_{t}=x_{t}+\gamma$ where $\gamma\sim U(-1.0,+1.0)$
  9. 9.
    Temporal warp: Stretch/compress: resample segment via $t^{\prime}=t_{\text{start}}+(t-t_{\text{start}})\cdot w$ where $w\sim U(0.8,1.2)$
  10. 10.
    Splice: Replace segment with values from different time: $\mathbf{x}^{\prime}_{t_{1}:t_{2}}=\mathbf{x}_{t_{3}:t_{3}+(t_{2}-t_{1})}$
  11. 11.
    Subtle drift: Very gentle linear trend: $x^{\prime}_{t}=x_{t}+\beta\cdot(t-t_{\text{start}})$ where $\beta\sim U(-0.002,+0.002)$
These simple transformations contrast sharply with test-time physical-space injections using multi-variant equations (exponential/sigmoid/polynomial drift, hydraulic-coupled ice effects, etc.).
E.0.2 Coverage Control Algorithm
To achieve target coverage $c_{\text{target}}$ (e.g., 10% for light tier, 20% for moderate tier), the injector uses iterative refinement:
  1. 1.
    Initialize: $n_{\text{segments}}\sim U(2,4)$, strength $\alpha=1.0$
  2. 2.
    For attempt = 1 to 3:
    1. (a)
      Generate $n_{\text{segments}}$ anomaly segments with strength $\alpha$
    2. (b)
      Compute realized coverage $c_{\text{realized}}$
    3. (c)
      If $|c_{\text{realized}}-c_{\text{target}}|<\delta$: return corrupted sequence
    4. (d)
      Else if $c_{\text{realized}}<c_{\text{target}}$: $\alpha\leftarrow\alpha\times U(1.1,1.4)$
    5. (e)
      Else: $\alpha\leftarrow\alpha\times U(0.7,0.9)$
  3. 3.
    Return best attempt (closest to target)
This adaptive mechanism maintains mean realized coverage 15.2% $\pm$ 3.1% across batches regardless of sequence characteristics.
Appendix F Test-Time Anomaly Injection
This appendix provides complete mathematical formulations for the 18 test anomaly types, each with 3-4 equation variants.
F.0.1 Example: Drift Anomaly (4 variants)
Variant 1 - Linear:
$Q^{\prime}(t)=Q(t)+\beta_{Q}\cdot(t-t_{\text{start}}),\quad H^{\prime}(t)=H(t)+\beta_{H}\cdot(t-t_{\text{start}})$ (45)
where $\beta_{Q}\sim U(-0.5,+0.5)$ ft3/s/hr, $\beta_{H}\sim U(-0.01,+0.01)$ ft/hr
Variant 2 - Exponential:
$Q^{\prime}(t)=Q(t)\cdot\exp[\alpha_{Q}\cdot(t-t_{\text{start}})],\quad H^{\prime}(t)=H(t)\cdot\exp[\alpha_{H}\cdot(t-t_{\text{start}})]$ (46)
where $\alpha_{Q}\sim U(-0.01,+0.01)$ hr-1, $\alpha_{H}\sim U(-0.005,+0.005)$ hr-1
Variant 3 - Sigmoid:
$Q^{\prime}(t)=Q(t)+\Delta Q\cdot\frac{1}{1+\exp[-k_{Q}(t-t_{\text{mid}})]},\quad H^{\prime}(t)=H(t)+\Delta H\cdot\frac{1}{1+\exp[-k_{H}(t-t_{\text{mid}})]}$ (47)
where $t_{\text{mid}}=(t_{\text{start}}+t_{\text{end}})/2$, $\Delta Q\sim U(-Q_{\text{mean}}/2,+Q_{\text{mean}}/2)$, $k_{Q}\sim U(0.1,0.5)$
Variant 4 - Polynomial:
$Q^{\prime}(t)=Q(t)+[a_{Q}(t-t_{\text{start}})^{2}+b_{Q}(t-t_{\text{start}})],\quad\text{similarly for }H$ (48)
where coefficients $a,b$ chosen to achieve target endpoint deviation $\sim U(0.1Q_{\text{mean}},0.3Q_{\text{mean}})$
F.0.2 Example: Ice Backwater (3 variants)
Variant 1 - Gradual onset:
$\alpha_{\text{ice}}(t)$ $=\alpha_{\max}\cdot\frac{t-t_{\text{start}}}{t_{\text{peak}}-t_{\text{start}}}\quad\text{for }t\in[t_{\text{start}},t_{\text{peak}}]$ (49)
$H^{\prime}(t)$ $=H(t)\cdot[1+\alpha_{\text{ice}}(t)]$ (50)
$Q^{\prime}(t)$ $=Q(t)\cdot[1-\beta_{\text{ice}}(t)]$ (51)
where $\alpha_{\max}\sim U(0.15,0.55)$, $\beta_{\max}\sim U(0,0.10)$, $t_{\text{peak}}-t_{\text{start}}\sim U(12,48)$ hours
Variant 2 - Abrupt onset, gradual recovery:
$\alpha_{\text{ice}}(t)=\begin{cases}\alpha_{\max}&t\in[t_{\text{start}},t_{\text{recover}}]\\ \alpha_{\max}\cdot\exp[-k(t-t_{\text{recover}})]&t>t_{\text{recover}}\end{cases}$ (52)
where $k\sim U(0.01,0.05)$ hr-1
Variant 3 - Periodic breakup events:
$\alpha_{\text{ice}}(t)=\alpha_{\text{base}}+\sum_{i=1}^{N_{\text{events}}}A_{i}\exp\left[-\frac{(t-t_{i})^{2}}{2\sigma_{i}^{2}}\right]$ (53)
where $\alpha_{\text{base}}\sim U(0.2,0.4)$, $N_{\text{events}}\sim U(2,5)$, event amplitudes $A_{i}\sim U(-0.3,-0.1)$ (negative for breakup), widths $\sigma_{i}\sim U(1,6)$ hours
Appendix G Expert Review Documentation
This appendix documents the structured expert review process for synthetic anomaly validation.
G.0.1 Review Protocol
Domain experts evaluated each anomaly type using standardized forms. Reviewers included:
  • A faculty member in geoscience with extensive experience in water resources and hydrological research
  • A research assistant from the University of Vermont Water Resources Institute with operational experience from a USGS internship, including exposure to real-world streamflow data quality challenges
Realism assessment (4-point scale):
  • 4 - Very realistic: Commonly observed in operational QA/QC workflows
  • 3 - Realistic: Plausible, occasionally observed
  • 2 - Questionable: Theoretically possible but rarely encountered
  • 1 - Unrealistic: Does not match field experience
Temporal scale validation: For each anomaly type, reviewers identified duration ranges appearing implausible relative to operational experience. For example, ice backwater durations <<24 hours were flagged as too brief for realistic ice formation/breakup dynamics.
Confound detection: Reviewers flagged injection regimes creating:
  • Artificially easy detection (e.g., spike magnitudes $10\sigma$ never observed in practice)
  • Ambiguous labeling (e.g., patterns indistinguishable from natural variability)
  • Physical impossibilities (e.g., rating curve violations)
G.0.2 Key Feedback Integration
Table 10 documents major revisions based on expert feedback.
Table 10: Expert feedback integration
Topic Initial Expert Feedback Change Made
Ice duration Macro: 72-144h Ice persists weeks-months in northern systems Extended to 72-520h
Backwater onset 50% instantaneous Backwater develops gradually as obstruction forms All use gradual onset (>>6h)
Spike magnitudes Some $8\sigma$ Never see such extremes in approved data Capped at $5\sigma$
Coverage tier 40% probability Reduce for challenging evaluation Reduced to 20%
Equation variants Some had 2 Need stronger separation for generalization claims Minimum 3 per type
Drift subtlety All >>0.1 ft3s-1day-1 Drift can be much more subtle Added subtle variant
G.0.3 Consensus Categorization
The final 18-type taxonomy reflects reviewer agreement on field-relevance:
High confidence (9 types): All reviewers rated 3-4 (realistic to very realistic), with specific operational examples provided:
  • Dropout: Battery failure, telemetry loss
  • Flatline: Frozen float, clogged intake
  • Ice backwater: Common Nov-Apr northern stations
  • Rating drift: Frequent after major floods, channel migration
Medium confidence (4 types): Mixed ratings (2-4), considered plausible but less frequently documented:
  • Time misalignment: Possible with manual data entry, rare in modern systems
  • Gate operation: Relevant for regulated sites, not universal
Low confidence (4 types): Majority rated 2 (questionable), retained for completeness:
  • Splice: Hard to imagine operational scenario
  • Warp: Never seen systematic time stretching
Debated (retained with caveat): Reviewers disagreed on anomaly vs. natural variability:
  • Noise burst: Could be natural turbulence or sensor issue, ambiguous
Appendix H Canadian Data Quality Filtering
This appendix details quality filtering applied to Canadian ECCC stations (Section 3.4.2).
H.0.1 Hydrologically-Motivated Checks
All checks apply to corrected series only, ensuring evaluation on windows where hydrologist-edited records behave as plausible physical time series.
1. Sufficient variability
$\text{CV}(H_{\text{corr}})=\frac{\sigma(H_{\text{corr}})}{\mu(H_{\text{corr}})}>0.10$ (54)
Rationale: Removes nearly flat records where water level hardly changes (e.g., failed sensors reporting constant values, tidal gauges at slack water). Anomaly detection is ill-posed on essentially constant series.
2. Monotonic rating relation
$\rho_{\text{Spearman}}(H_{\text{corr}},Q_{\text{corr}})>0.5$ (55)
Rationale: Enforces physically plausible behavior where larger stages generally correspond to larger discharges. Violations suggest either: (a) severe data quality issues making the window unsuitable for evaluation, or (b) complex hydraulics (backwater, tidal influence) where the station may not be appropriate for standard QA/QC methods.
3. Reasonable rating curve exponent
Fit $\log Q=\log a+b\log H$ using robust regression (RANSAC), require:
$b\in[0.5,10]$ (56)
Rationale: Natural open-channel flow typically exhibits $b\in[1.5,3.0]$ based on hydraulic geometry [chow1988applied]. We relax to [0.5, 10] to accommodate weirs ($b\approx 1.5$), rectangular channels ($b\approx 1.0$), and complex cross-sections, while excluding physically implausible extremes.
4. Moderate rating curve fit
$R^{2}\geq 0.3$ (57)
Rationale: Deliberately modest threshold allows hydraulic complexity (hysteresis, backwater) while removing windows where stage explains essentially none of discharge variability.
5. Sufficient valid data fraction
$\frac{\#\{t:Q_{\text{corr},t}>0\text{ and }H_{\text{corr},t}>0\}}{\#\{\text{all timesteps}\}}\geq 0.70$ (58)
Rationale: Ensures enough usable information for robust anomaly detection features. Windows with $30\%$ missing/invalid data lack statistical power.
6. Limited flatline segments
Compute consecutive differences: $\Delta H_{t}=|H_{\text{corr},t}-H_{\text{corr},t-1}|$. Require:
$\frac{\#\{t:\Delta H_{t}<0.001\text{ ft}\}}{\#\{\text{all timesteps}\}}<0.30$ (59)
Rationale: Natural hydrographs exhibit continuous variation. If $30\%$ of consecutive values are identical (within 0.001 ft precision), the gauge is likely frozen or malfunctioning. This preserves low-variability baseflow periods while removing structural sensor failures.
H.0.2 Station-Level Filtering
After temporal alignment of the four series (stage raw/corrected, discharge raw/corrected), we compute missing data fraction:
$f_{\text{missing}}=\frac{\#\{\text{timesteps missing any of 4 series}\}}{\#\{\text{all timesteps}\}}$ (60)
Stations with $f_{\text{missing}}>0.05$ (losing $5\%$ of hourly timesteps) are excluded entirely from the Canadian test set.
Rationale: High missing data rates indicate either: (a) inconsistent data collection between raw and corrected archives, or (b) stations undergoing major operational changes. Both scenarios compromise evaluation validity.
H.0.3 Filtering Statistics
Across the full ECCC archive:
  • Initial stations with raw + corrected data: 1,847
  • After station-level filtering ($5\%$ missing): 1,203 (65.1% retention)
  • After window-level filtering (6 criteria): 487 stations with eligible windows (26.4% overall retention)
  • Random sampling for evaluation: 100 stations
The conservative filtering ensures evaluation on high-quality, hydrologically meaningful windows where model performance reflects genuine anomaly detection capability rather than artifacts of poor data quality.